### v2

# Most Probable KAM Tori in Stochastic Hamiltonian Systems

Abstract
This paper conducts an in-depth investigation into the impact of stochastic perturbations-particularly multiplicative noise-on the integrable structures of Hamiltonian systems, with a central focus on developing a KAM theory for stochastic Hamiltonian dynamics. We begin by deriving the Onsager-Machlup functional for Hamiltonian systems driven by multiplicative noise and identifying the most probable transition paths of system trajectories. This analysis reveals the fundamental differences in how additive versus multiplicative noise influences the integrability of Hamiltonian systems. Building upon this, we establish a large deviation principle for the system and derive a rate function that quantitatively characterizes trajectory deviations, especially in the regime of rare events. The main contribution of this work lies in demonstrating that, under the small noise limit, the quasi-periodic invariant tori of the unperturbed system persist in a probabilistic sense, indicating the stability of KAM structures under stochastic perturbations. Furthermore, we show that the exponential rate of deviation from the invariant tori exactly matches the large deviation rate function, thus providing a quantitative characterization of the structural persistence and fluctuation geometry of quasi-periodic motions in stochastic Hamiltonian systems. These results extend the classical KAM framework to a stochastic setting and offer new insights into the behavior of complex dynamical systems under the influence of noise.
70H08,
60H10,
82C35,
60F10,
70L05,
Stochastic Hamiltonian systems,
Invariant tori,
Onsager-Machlup functional,
Large deviation,
KAM theory,
keywords:
[class=MSC]
keywords:
[…]
, thanks: corresponding author..
Statements and Declarations
  • Ethical approval
    Not applicable.
  • The Data Availability Statement.
    No datasets were generated or analysed during the current study.
  • The Conflict of Interest Statement.
    We have no conflicts of interest to disclose.
  • Funding
    The second author (Y. Li) was supported by National Natural Science Foundation of China (Grant No. 12071175, 12471183, 12531009).
1 Introduction
This paper focuses on the persistence of invariant tori in stochastic Hamiltonian systems, particularly examining their stability under random perturbations. We consider the following stochastic Hamiltonian system driven by multiplicative noise:
$\begin{cases}\,{\rm d}\theta(t)=\frac{\partial H}{\partial I}(\theta,I)\,{\rm d}t+\sigma_{\theta}(t,\theta,I)\,{\rm d}W_{\theta}(t),\\ \,{\rm d}I(t)=-\frac{\partial H}{\partial\theta}(\theta,I)\,{\rm d}t+\sigma_{I}(t,\theta,I)\,{\rm d}W_{I}(t),\end{cases}$ (1)
where $H(\theta,I)$ is the Hamiltonian, $\theta$ represents the angle variables, and $I$ represents the action variables. $\sigma_{\theta}(t,\theta,I)$ and $\sigma_{I}(t,\theta,I)$ represent the strengths of the random perturbations, and $W_{\theta}(t)$ and $W_{I}(t)$ are standard Wiener processes. The corresponding symplectic form of this system is given by
$\mathrm{d}X(t)=J\nabla H(X(t))\,\mathrm{d}t+\sigma(t,X(t))\,\mathrm{d}W(t),$ (2)
where $J$ is the standard symplectic matrix defined by
$J:=\begin{pmatrix}0&I_{n}\\ -I_{n}&0\end{pmatrix}\in\mathbb{R}^{2n\times 2n},$
and $\nabla H(X)$ denotes the gradient of the Hamiltonian with respect to $X=(\theta,I)$, given by
$\nabla H(X)=\begin{pmatrix}\frac{\partial H}{\partial\theta}\\ \frac{\partial H}{\partial I}\end{pmatrix}$,
and the diffusion matrix $\sigma(t,X)$ is defined as the block-diagonal matrix
$\sigma(t,X):=\begin{pmatrix}\sigma_{\theta}(t,\theta,I)&0\\ 0&\sigma_{I}(t,\theta,I)\end{pmatrix}\in\mathbb{R}^{2n\times 2n}.$
By integrating the Onsager-Machlup functional, large deviation principles, and KAM theory, we establish a framework for analyzing the most probable paths and stability of the system under stochastic conditions, thereby revealing the persistence of invariant tori in the most probable context.
Hamiltonian systems are foundational in classical mechanics, with applications across physics, astronomy, mechanical systems, and more. The core framework, dating back to the early 19th century, was introduced by William Rowan Hamilton. Hamiltonian mechanics describes a system’s state via generalized coordinates $\theta$ and conjugate momenta $I$, while the Hamiltonian function $H(\theta,I)$ represents the system’s total energy, encompassing both kinetic and potential energy. The evolution of such a system is governed by the Hamiltonian equations
$\mathrm{d}X(t)=J\nabla H(X(t))\,\mathrm{d}t.$
This formulation not only provides a precise description of system dynamics but also ensures energy conservation and preservation of symplectic geometry. However, when subjected to perturbations, the system’s behavior can become complex. In particular, understanding how to maintain long-term stability under small perturbations presents a critical challenge.
To address these challenges, the KAM theory was a significant breakthrough in the 20th century. Kolmogorov [25] hypothesized that when Hamiltonian systems are subject to small perturbations, some of the invariant tori (regular orbital structures) would persist and avoid chaotic behavior. Arnold [1] and Moser [34] subsequently provided rigorous proofs of this conjecture, formalizing what is now known as the KAM theory. The core result demonstrates that under small perturbations, most of invariant tori continue to exist, preserving the system’s quasi-periodic motions. This result profoundly advanced the study of Hamiltonian systems’ stability under perturbations. Subsequent relevant developments are referred to as [17, 37, 26, 36, 31, 6, 9, 38], and so on.
The original KAM framework was designed primarily for deterministic perturbations, leaving the question of its applicability in stochastic settings unresolved. As a result, validating the extension of the KAM theory under stochastic perturbations and quantifying the stability of invariant tori has emerged as a central problem in stochastic Hamiltonian system research. Recent studies have made significant progress in this area. Wu [46] established a framework for both large and moderate deviations in stochastic Hamiltonian systems, providing a quantitative assessment of the probability of rare events. Talay [44] explored how stochastic Hamiltonian systems asymptotically converge to an invariant measure, with a focus on the exponential nature of this convergence. Lázaro-Camí and Ortega [27] investigated the impact of stochastic noise on classical Hamiltonian dynamics, elucidating how structural preservation can be described in stochastic environments. Li [30] proposed an averaging principle framework for completely integrable stochastic Hamiltonian systems, simplifying the analysis of their long-term behavior under small stochastic perturbations. Zhang [47] examined stochastic flows within Hamiltonian systems and introduced new computational methods based on the Bismut formula. Some of the latest research references are [29, 11, 8] and so on.
The value of this paper lies in the development of a novel analytical framework for investigating the stability of invariant tori in stochastic Hamiltonian systems. By integrating the Onsager-Machlup functional, large deviation theory, and KAM theory, we provide new insights into the behavior of such systems under stochastic perturbations, particularly in the presence of multiplicative noise. The main contributions of this work are summarized as follows:
  1. 1.
    We establish, for the first time, a rigorous formulation of the Onsager-Machlup functional for stochastic Hamiltonian systems with multiplicative noise, thereby extending previous results which were restricted to additive noise. Our formulation accommodates general nonlinear diffusion coefficients that depend on the state variables and captures the non-Gaussian nature of path distributions. To overcome technical challenges arising from non-adapted path perturbations, we further develop a conditional expectation comparison principle and a small ball probability estimate tailored for the multiplicative noise setting.
  2. 2.
    Using the Ansatz method, we rigorously prove that the most probable path satisfies a modified deterministic Hamiltonian equation. On this basis, we derive an explicit large deviation principle and the corresponding rate functional via the Freidlin–Wentzell framework, providing a quantitative characterization of the probability that rare trajectories deviate from the invariant tori.
  3. 3.
    We extend the classical KAM theory to the stochastic setting, and prove that under the most probable path interpretation, invariant tori of nearly integrable Hamiltonian systems remain stable under small stochastic perturbations, provided the non-degeneracy and Diophantine conditions are satisfied. This result offers a new probabilistic perspective on the long-term stability of Hamiltonian systems.
The Onsager-Machlup functional is a key analytical tool for studying path probabilities and rare events. It identifies the most probable transition paths among all smooth paths in a noise-driven system. This functional quantifies the likelihood of different paths in a probabilistic setting, playing a role similar to the action functional in classical mechanics. The Onsager-Machlup functional originated from the work of Onsager [35] and Machlup [32] in 1953, where it was introduced to describe the probability density functional for diffusion processes with linear drift and constant diffusion coefficients. Subsequently, in 1957, Tisza and Manning [45] extended its application to nonlinear equations, and in the same year, Stratonovich [42] provided a rigorous theoretical framework. Recent developments refer to [33, 3, 5, 28, 4], and so on. In stochastic Hamiltonian systems, the Onsager-Machlup functional is a powerful tool for identifying the most probable path under random perturbations, and it provides a foundation for further large deviation analysis.
The origins of large deviation theory and its associated research can be traced back to the early 20th century. Cramér [10] and Sanov [40] made foundational contributions to the study of large deviations in sequences of independent and identically distributed random variables. Subsequently, Donsker and Varadhan [14, 15, 16] systematically studied large deviations in the context of Markov processes and their connection to ergodicity. Their work introduced key concepts such as Varadhan’s integral lemma and the contraction principle, which are not only central results in large deviation theory but also established deep connections with other fields of mathematics (see [43, 12, 13, 18]). In the 1970s, Freidlin and Wentzell [19] extended the theory by applying it to stochastic dynamical systems and stochastic differential equations, particularly in the context of small perturbations. The Freidlin-Wentzell framework describes the probability of a system deviating from its most likely path and introduces the rate function to quantify the distribution of deviations from typical behavior. In our consideration, this rate function, derived from extremal analysis of the Onsager-Machlup functional, allows for effective prediction of the probability distribution of deviations from invariant tori. This provides a novel perspective for quantifying system stability under stochastic perturbations and is of significant importance for understanding the long-term stability of stochastic Hamiltonian systems.
A major difficulty in this work arises from the more challenging multiplicative noise setting, in which the diffusion coefficient $\sigma(t,X_{t})$ depends nonlinearly on the current state $X_{t}$. This leads to non-Gaussian path distributions and introduces significant analytical challenges. Specifically, the classical Cameron–Martin transformation no longer preserves the law of the process, making small ball estimates and conditional expectations difficult to control using standard Gaussian techniques. Moreover, path perturbations no longer satisfy linear superposition, rendering traditional tools such as orthogonal expansions and characteristic function methods inapplicable. The lack of explicit density formulas for non-Gaussian paths further complicates the analysis of conditional expectations and path-dependent functionals.
Notably, foundational results in Onsager-Machlup theory—such as Theorem 2.12 and Theorem 2.14—fail to hold under such general settings. To address these challenges, we first establish a small ball probability estimate (Lemma 2.19), which serves as a key technical tool. Building on this result, we introduce a conditional expectation comparison principle, which extends Theorems 2.12 and 2.14, leading to Lemmas 2.20 and 2.21 that are tailored for the multiplicative noise framework.
On this foundation, we rigorously characterize the Onsager-Machlup functional for the multiplicative noise system using advanced non-Gaussian techniques, including the Radon-Nikodym derivative structure under Girsanov’s transformation, non-adapted Skorohod integration, small ball probability estimates, and conditional expectation bounds. We show that the Onsager-Machlup functional is given by
$OM(\varphi,\dot{\varphi})=\int_{0}^{1}\left\|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right\|^{2}\,\mathrm{d}s-\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,\mathrm{d}s.$
Next, we apply the Ansatz method and the variational approach to minimize the Onsager-Machlup functional and determine the most probable continuous path of the system. We prove that this most probable path corresponds to the solution of the modified deterministic Hamiltonian equation
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}.$
By combining the Onsager-Machlup functional with the Freidlin-Wentzell large deviation theory [28], we establish a large deviation principle for the stochastic Hamiltonian system. Furthermore, we derive the associated rate function
$S(\varphi)=\frac{1}{2}\int_{0}^{1}\left\|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right\|^{2}\,\mathrm{d}s,$
which quantifies the deviation of rare paths from the most probable trajectory. We observe that this rate function is closely related to the Onsager-Machlup functional, with the trace-related term in the latter corresponding to higher-order contributions in the small noise limit and hence vanishing under the large deviation scaling.
Finally, by employing the classical KAM theory, we prove that for a nearly integrable Hamiltonian system, the invariant tori remain stable in the most probable sense under small stochastic perturbations, provided the non-degeneracy and Diophantine conditions are satisfied. Notably, we find that the probability of deviation from these invariant tori is precisely characterized by the rate function derived above.
The structure of this paper is as follows: In Section 2, we review some basic definitions of spaces and norms, introduce the concept of the Onsager-Machlup functional, and present several key technical lemmas. In Section 3, we derive the Onsager-Machlup functional for stochastic Hamiltonian systems. In Section 4, we prove that the most probable path of a stochastic Hamiltonian system corresponds to the stable solution of its associated deterministic Hamiltonian system. A specific example of a one-dimensional stochastic harmonic oscillator is provided to illustrate our results. In Section 5, we derive the large deviation principle for stochastic Hamiltonian systems. Finally, in Section 6, we extend the preceding results to the case of nearly integrable stochastic Hamiltonian systems, proving a stochastic version of the KAM theory and providing specific examples to illustrate our findings.
2 Preliminaries
2.1 Basic space and norm
In this section, we review some fundamental definitions and results concerning asymptotic limits in the Wiener space (see reference [44]), as well as several basic definitions of functional spaces and norms (see reference [20, 22]).
Let $W=\left\{W_{t},~t\in[0,1]\right\}$ be a Brownian motion (Wiener process) defined in the complete filtered probability space $(\Omega,\mathcal{F},\left\{\mathcal{F}_{t}\right\}_{t\geq 0},\mathbb{P})$. Here, $\Omega$ represents the space of continuous functions vanishing at zero, and $\mathbb{P}$ denotes the Wiener measure. Let $\mathbb{H}:=L^{2}([0,1],\mathbb{R}^{n})$ be a Hilbert space and $\mathbb{H}^{1}$ be the Cameron-Martin space defined as follows:
$\mathbb{H}^{1}$ $=\left\{f:[0,1]\to\mathbb{R}^{n}\in\mathbb{H}^{1}~\big{|}~f(0)=0,f~\text{is absolutely continuous functions and}~f^{\prime}\in\mathbb{H}\right\}.$
The scalar product in $\mathbb{H}^{1}$ is defined as follows:
$\langle f,g\rangle_{\mathbb{H}^{1}}=\langle f^{\prime},g^{\prime}\rangle_{\mathbb{H}}$
for all $f,g\in\mathbb{H}^{1}$. Let $\mathcal{P}:\mathbb{H}^{1}\to\mathbb{H}^{1}$ be an orthogonal projection with $dim\mathcal{P}\mathbb{H}^{1}<\infty$ and the specific expression
$\mathcal{P}h=\sum_{i=1}^{n}\langle h_{i},f\rangle h_{i},$
where $(h_{1},...,h_{n})$ is a set of orthonormal basis in $\mathcal{P}\mathbb{H}^{1}$. In addition, we can also define the $\mathbb{H}^{1}$-valued random variable
$\mathcal{P}^{W}=\sum_{i=1}^{n}\bigg{(}\int_{0}^{1}{h_{i}^{\prime}}\,{\rm d}W_{s}\bigg{)}h_{i},$
where $\mathcal{P}^{W}$ is independent of the choice of the orthonormal basis $(h_{1},\ldots,h_{n})$.
Definition 2.1.
We say that a sequence of orthogonal projections $\mathcal{P}_{n}$ on $\mathbb{H}^{1}$ is an approximating sequence of projections, if $dim\mathcal{P}_{n}\mathbb{H}^{1}<\infty$ and $\mathcal{P}_{n}$ converges strongly to the identity operator $I$ in $\mathbb{H}^{1}$ as $n\to\infty$.
Definition 2.2.
We say that a semi-norm $\mathcal{N}$ on $\mathbb{H}^{1}$ is measurable, if there exists a random variable $\tilde{\mathcal{N}}$, satisfying $\tilde{\mathcal{N}}$$\infty$ a.s, such that for any approximating sequence of projections $\mathcal{P}_{n}$ on $\mathbb{H}^{1}$, the sequence $\mathcal{N}(\mathcal{P}^{W}_{n})$ converges to $\tilde{\mathcal{N}}$ in probability and $\mathbb{P}(\tilde{\mathcal{N}}\leq\epsilon)>0$ for any $\epsilon>0$. Moreover, if $\mathcal{N}$ is a norm on $\mathbb{H}^{1}$, then we call it a measurable norm.
Definition 2.3.
We define the space $C^{2}_{b}([0,1],\mathbb{R}^{n\times n})$ as the set of matrix-valued functions $g(t)$ that satisfy the following conditions:
  1. 1.
    For every $t\in[0,1]$, $g(t)$ is a real symmetric matrix;
  2. 2.
    $g(t)$ is twice continuously differentiable;
  3. 3.
    $g(t)$ is bounded, meaning there exists a constant $M>0$ such that $\|g(t)\|_{2}\leq M$ for all $t\in[0,1]$, where $\|g(t)\|_{2}$ denotes the spectral norm of the matrix $g(t)$.
Definition 2.4.
We define the space $C^{2}_{b}([0,1]\times\mathbb{R}^{d},\mathbb{R}^{n\times n})$ as the set of matrix-valued functions $g(t,x)$ that satisfy the following conditions:
  1. 1.
    For every $(t,x)\in[0,1]\times\mathbb{R}^{d}$, $g(t,x)$ is a real symmetric matrix;
  2. 2.
    $g(t,x)$ is twice continuously differentiable with respect to all variables $(t,x)$;
  3. 3.
    $g(t,x)$ is bounded, meaning there exists a constant $M>0$ such that $\|g(t,x)\|_{2}\leq M$ for all $(t,x)\in[0,1]\times\mathbb{R}^{d}$, where $\|g(t,x)\|_{2}$ denotes the spectral norm of the matrix $g(t,x)$.
Definition 2.5.
For a matrix $A=[a_{ij}]\in\mathbb{R}^{n\times n}$, its spectral norm is defined as:
$\|A\|_{2}=\max_{\|x\|_{2}=1}\|Ax\|_{2}.$
Alternatively, it can be expressed as:
$\|A\|_{2}=\sigma_{\text{max}}(A),$
where $\sigma_{\text{max}}(A)$ denotes the largest singular value of the matrix $A$. The singular values of $A$ are the square roots of the eigenvalues of the matrix $A^{T}A$.
Definition 2.6.
Let $A=[a_{ij}]\in\mathbb{R}^{n\times n}$. Its Hilbert–Schmidt norm (also known as the Frobenius norm) is given by
$\|A\|_{\mathrm{HS}}\;=\;\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{n}\bigl{\lvert}a_{ij}\bigr{\rvert}^{2}}\;=\;\sqrt{\mathrm{trace}\bigl{(}A^{T}A\bigr{)}}.$
Equivalently, if $\{\sigma_{k}(A)\}$ are the singular values of $A$, then
$\|A\|_{\mathrm{HS}}\;=\;\sqrt{\sum_{k=1}^{n}\sigma_{k}(A)^{2}}.$
Thus the Hilbert–Schmidt norm captures the sum of the squares of all entries (or, equivalently, the square root of the sum of the squares of all singular values) of the matrix.
Definition 2.7.
Let $f$ be a function defined on a compact domain $\Omega\subset\mathbb{R}^{d}$. The supremum norm (also known as the maximum norm) of $f$ on $\Omega$ is defined by
$\|f\|_{0;\Omega}:=\sup_{x\in\Omega}|f(x)|,$
where $|\cdot|$ denotes the Euclidean norm in $\mathbb{R}^{d}$. This norm quantifies the largest absolute value attained by $f$ over the domain $\Omega$.
2.2 Onsager-Machlup functional
In the problem of finding the most probable path of a diffusion process, the probability of a single path is zero. Instead, we can search for the probability that the path lies within a certain region, which could be a tube along a differentiable function. This tube is defined as
$K(\varphi,\epsilon)=\{x-x_{0}\in\mathbb{H}^{1}\mid\varphi-x_{0}\in\mathbb{H}^{1},\|x-\varphi\|\leq\epsilon,\epsilon>0\}.$
Once $\epsilon>0$ is given, the probability of the tube can be expressed as
$\mu_{x}(K(\varphi,\epsilon))=P\left(\{\omega\in\Omega\mid X_{t}(\omega)\in K(\varphi,\epsilon)\}\right),$
allowing us to compare the probabilities of the tubes for all $\varphi-x_{0}\in\mathbb{H}^{1}$.
Thus, the Onsager-Machlup function can be defined as the Lagrangian function that gives the most probable tube. We now introduce the definitions of the Onsager-Machlup function and the Onsager-Machlup functional.
Definition 2.8.
Consider a tube surrounding a reference path $\varphi_{t}$ with initial value $\varphi_{0}=x$ and $\varphi_{t}-x$ belongs to $\mathbb{H}^{1}$. Assuming $\epsilon$ is given and small enough, we estimate the probability that the solution process $X_{t}$ is located in that tube as
$\mathbb{P}\left\{\|X-\varphi\|\leq\epsilon\right\}\propto C(\epsilon){\rm exp}\left\{-\frac{1}{2}\int_{0}^{1}{OM(t,\varphi,\dot{\varphi})}\,{\rm d}t\right\},$
where $\propto$ denotes the equivalence relation for $\epsilon$ small enough. We call the integrand $OM(t,\varphi,\dot{\varphi})$ the Onsager-Machulup function and also call integral $\int_{0}^{1}{OM(t,\varphi,\dot{\varphi})}\,{\rm d}t$ the Onsager-Machulup functional. In analogy to classical mechanics, we also refer to the Onsager-Machulup function as the Lagrangian function and the Onsager-Machulup functional as the action functional.
2.3 KAM Theory
In Hamiltonian mechanics, invariant tori describe the set of solutions exhibiting quasi-periodic motion. These tori are higher-dimensional analogs of closed orbits, which arise when the system evolves with incommensurate frequencies. Systems with invariant tori are often referred to as integrable systems, as their dynamics are regular, confined to these tori, and predictable in phase space.
KAM theory focuses on investigating the stability of these invariant tori under small perturbations. For a nearly integrable Hamiltonian system, where the Hamiltonian is composed of an integrable part plus a small perturbation term, KAM theory asserts that as long as the perturbation is sufficiently small and certain conditions are met, most of the original invariant tori persist, albeit with slight deformations.
We will use the following theorem, which is stated as Theorem 3.1 in [26].
Theorem 2.9 (Koudjinan).
Consider a Hamiltonian of the form $H(I,\theta)=K(I)+P(I,\theta)$, where $I\in\mathcal{D}\subset\mathbb{R}^{d}$ are the action variables and $\theta\in\mathbb{T}^{d}$ are the angle variables. Here, $K(I)$ and $P(I,\theta)$ are $C^{l}$-smooth functions with $K,P\in C^{l}(\mathcal{D}\times\mathbb{T}^{d})$, where $\mathcal{D}$ is a non-empty bounded domain in $\mathbb{R}^{d}$. If $K$ is non-degenerate and $l>2\nu>2d$, then all the KAM tori of the integrable system $K$ whose frequency are $(\alpha,\tau)$-Diophantine, with $\alpha\simeq\epsilon^{1/2-\nu/l}$ and $\tau:=\nu-1$, do survive, being only slightly deformed, where $\epsilon$ is the $C^{l}$-norm of the perturbation $P$. Moreover, letting $\mathcal{K}$ be the corresponding family of KAM tori of $H$, we have
$\mathrm{meas}(\mathcal{D}\times\mathbb{T}^{d}\setminus\mathcal{K})=O(\epsilon^{1/2-\nu/l}).$
This theorem provides a more refined theoretical foundation for the persistence of invariant tori in finitely differentiable Hamiltonian systems, extending the classical KAM theory to the case where the Hamiltonian is only finitely smooth. It demonstrates that, even under conditions of finite differentiability, a significant portion of the invariant tori remains stable. This stability implies that, despite perturbations, many quasi-periodic motions can still exist and maintain their regularity in phase space. This result enhances the robustness of the KAM theory, showing that the structure of Hamiltonian systems can exhibit notable stability even under less stringent smoothness conditions.
2.4 Technical theorems
In this section, we will introduce several commonly utilized technical theorems. Throughout this paper, unless otherwise specified, the norm $\|\cdot\|$ refers to the supremum norm $\|\cdot\|_{0;\Omega}$. We use $C$ to denote a generic constant, which may vary from line to line. The symbol $n$ denotes the spatial dimension, which is assumed to be an arbitrary but finite positive integer. The notation $\mathbb{E}(A\mid B)$ stands for the conditional expectation of the random variable $A$ given the event or $\sigma$-algebra $B$.
When we derive the Onsage-Machup functional of stochastic differential equations with additive noise, the following theorem is the most basic one, as it ensures that we handle each term separately. For a complete proof, see [23], pages 536-537.
Theorem 2.10.
For a fixed integer $N\geq 1$, let $X_{1},...,X_{N}\in\mathbb{R}$ be $N$ random variables defined on $(\Omega,\mathcal{F},\left\{\mathcal{F}_{t}\right\}_{t\geq 0},\mathbb{P})$ and $\left\{D_{\epsilon};\epsilon>0\right\}$ be a family of sets in $\mathcal{F}$. Suppose that for any $c\in\mathbb{R}$ and any i=1,,Ni=1,...,N, we have
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cX_{i}\right\}\big{|}D_{\epsilon}\right)\leq 1.$
Then
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{\sum_{i=1}^{N}cX_{i}\right\}\big{|}D_{\epsilon}\right)=1.$
The following theorem originates from the seminal work of Shepp and Zeitouni [41], where they investigated the conditional behavior of exponential moments of Itô integrals under small-ball constraints (see page 2 of their paper). In particular, they established that for any deterministic integrand in $L^{2}[0,1]$, the conditional exponential moment approaches unity as the Brownian path is conditioned to remain within an increasingly small neighborhood. We formally restate their result below:
Theorem 2.11.
Let $f\in L^{2}[0,1]$ be a deterministic function. Then, for any “reasonable” norm $\|\cdot\|$ on $C_{0}[0,1]$ (for example, the maximum norm or the Hölder norm), we have
$\lim\limits_{\epsilon\to 0}\mathbb{E}\left(\exp\left\{\int_{0}^{1}f(s)\,\mathrm{d}W_{s}\right\}\,\middle|\,\|W\|<\varepsilon\right)=1,$
where $W_{t}$ denotes a standard Brownian motion.
To extend this framework to a more general class of norms and vector-valued integrands, we adopt the concept of measurable norms on the Cameron–Martin space $\mathbb{H}^{1}$. The following two theorems, stated as Theorems 8 and 10 in [21] and due to Gilles, provide such a generalization.
Theorem 2.12.
Let $\mathcal{N}$ be a measurable norm on $\mathbb{H}^{1}$. For any deterministic function $f\in L^{2}([0,1],\mathbb{R}^{n})$, we have
$\lim\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{{\int_{0}^{1}\left\langle{f(s)},\,{\rm d}W_{s}\right\rangle}\right\}\big{|}\tilde{\mathcal{N}}\leq\epsilon\right)=1,$
where $W_{t}$ denotes an $n$-dimensional standard Brownian motion with independent components and $\tilde{\mathcal{N}}$ is defined by Definition 2.2.
Definition 2.13.
We say that an operator $\mathcal{S}:\mathbb{H}\to\mathbb{H}$ is nuclear, if
$\sum_{n=1}^{\infty}\left|\langle\mathcal{S}e_{n},k_{n}\rangle\right|<\infty$
for any orthonormal sequences $B_{1}=(e_{n})_{n}$ and $B_{2}=(k_{n})_{n}$ in $\mathbb{H}$.
We define the trace of a nuclear operator $\mathcal{S}$ as follows:
$\mathrm{Tr}(\mathcal{S})=\sum_{n=1}^{\infty}\langle\mathcal{S}e_{n},e_{n}\rangle$
for any orthonormal sequence $B=(e_{n})_{n}$ in $\mathbb{H}$. The definition of trace is independent of the orthonormal sequences we choose. For a given symmetric function $f\in L^{2}([0,1]^{2})$, the Hilbert-Schmidt operator $\mathcal{S}(f):\mathbb{H}\to\mathbb{H}$ is defined as
$\left(\mathcal{S}\left(f\right)\right)(h)(t)=\int_{0}^{t}{f(t,s)h(s)}\,{\rm d}s$
is nuclear if $\sum_{n=1}^{\infty}\langle\mathcal{S}e_{n},e_{n}\rangle<\infty$ for any orthonormal sequence $B=(e_{n})_{n}$ in $\mathbb{H}$. When the function $f$ is continuous and the operator $\mathcal{S}(f)$ is nuclear, the trace of $f$ has the following expression(see [2]):
$\mathrm{Tr}\left(f\right):=\mathrm{Tr}\left(\mathcal{S}(f)\right)=\int_{0}^{1}{f(s,s)}\,{\rm d}s.$
Furthermore, when f(s,t){f(s,t)} is a continuous $n\times n$ covariance kernel in the square $0\leq s,t\leq 1$, the corresponding operator $\mathcal{S}$ is nuclear and the expression for its trace is as follows:
$\mathrm{Tr}\left(f\right)=\mathrm{Tr}\left(\mathcal{S}(f)\right)=\int_{0}^{1}{\mathrm{Tr}\left(f(s,s)\right)}\,{\rm d}s.$
Theorem 2.14.
Let $f$ be a symmetric function in $L^{2}([0,1]^{2})$ and let $\mathcal{N}$ be a measurable norm. If $\mathcal{S}(f)$ is nuclear, then
$\lim\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{\int_{0}^{1}{\int_{0}^{1}{f(s,t)}\,{\rm d}W_{s}}\,{\rm d}W_{t}\right\}\big{|}\tilde{\mathcal{N}}\leq\epsilon\right)=e^{-\mathrm{Tr}(f)}.$
The following theorem provides a classical estimate for the small ball probability of Brownian motion, which serves as a foundational result for the developments in this paper. A detailed explanation can be found on page 2 of [29].
Theorem 2.15.
Let $W(t)$ be a standard Brownian motion on $\mathbb{R}$, and define the first exit time
$\tau:=\inf\{s>0:|W(s)|\geq 1\}.$
Then for any $\varepsilon>0$, the small ball probability under the supremum norm satisfies the following scaling identity:
$\mathbb{P}\left(\sup_{0\leq t\leq 1}|W(t)|\leq\varepsilon\right)=\mathbb{P}\left(\sup_{0\leq t\leq\varepsilon^{-2}}|W(t)|\leq 1\right)=\mathbb{P}\left(\tau\geq\varepsilon^{-2}\right).$
Moreover, as $\varepsilon\to 0$, the probability exhibits the following logarithmic asymptotic behavior:
$\log\mathbb{P}\left(\sup_{0\leq t\leq 1}|W(t)|\leq\varepsilon\right)\sim-\frac{\pi^{2}}{8\varepsilon^{2}}.$
Remark 1.
heorem 2.15 provides a classical small ball estimate for one-dimensional Brownian motion. Similarly, by applying the multiplication principle for independent events, an analogous small ball estimate can be derived for $n$-dimensional Brownian motion. Although the constant may depend on the dimension $n$, since $n$ is finite, this dependence can be absorbed into the constant $C$, yielding the same asymptotic rate.
Below is the general theorem of the KAM scheme; for a comprehensive proof, refer to Appendix B of [26].
Theorem 2.16.
Let $r>0$, $0<\bar{\sigma}\leq 1$, $0<2\sigma<s\leq 1$, and $\mathscr{D}\subset\mathbb{R}^{d}$ be a non-empty, bounded domain. Consider the Hamiltonian
$H(I,\theta):=K(I)+P(I,\theta),$
where $K,P\in\mathcal{A}_{r,s}(\mathscr{D})$. Assume the following conditions hold:
$\det K_{II}(I)$ $\neq 0,$ $T(I)$ $=K_{II}(I)^{-1},$ $\forall I\in\mathscr{D},$ (3)
$\|K_{II}\|_{r,\mathscr{D}}$ $\leq\mathrm{K},$ $\|T\|_{\mathscr{D}}$ $\leq\mathrm{T},$
$\|P\|_{r,s,\mathscr{D}}$ $\leq\varepsilon,$ $K_{I}(\mathscr{D})$ $\subset\Delta_{\alpha}^{\tau}.$
Define the parameters
$\theta$ $=\mathrm{T}\mathrm{K},$ $\lambda$ $=\log\rho^{-1},$ $\kappa$ $=6\sigma^{-1}\lambda,$ (4)
$\tilde{r}$ $\leq\frac{r}{32d\theta},$ $\overline{r}$ $\leq\min\left\{\frac{\alpha}{2d\mathrm{K}\kappa^{\nu}},\tilde{r}\right\}$, $\breve{r}$ $=\frac{\tilde{r}\bar{\sigma}}{16d\theta},$
$\overline{s}$ $=s-\frac{2}{3}\sigma,$ $s^{\prime}$ $=s-\sigma,$ $\mathrm{L}$ $=C_{0}\frac{\theta^{2}\varepsilon}{r\breve{r}}.$
Furthermore, assume
$\sigma^{-\nu}\frac{\varepsilon}{\alpha r}\leq\rho\leq\frac{1}{4},\qquad r\leq\frac{\alpha}{\mathrm{K}}\sigma^{\nu},\qquad\mathrm{L}\leq\frac{\bar{\sigma}}{3}.$ (5)
Then, there exists a diffeomorphism $G:D_{\overline{r}}(\mathscr{D})\to G(D_{\overline{r}}(\mathscr{D}))$ and a symplectic change of coordinates $\phi^{\prime}=\text{id}+\tilde{\phi}:D_{\overline{r}/2,s^{\prime}}(\mathscr{D}^{\prime})\to D_{\overline{r}+r\sigma/3,\overline{s}}(\mathscr{D})$, where $\mathscr{D}^{\prime}:=G(\mathscr{D})$, such that
$\begin{cases}H\circ\phi^{\prime}=H^{\prime}:=K^{\prime}+P^{\prime},\\ \partial_{I^{\prime}}K^{\prime}\circ G=\partial_{I}K,\quad\det\partial_{I^{\prime}}^{2}K^{\prime}\circ G\neq 0\quad\text{on }\mathscr{D},\end{cases}$ (6)
where $K^{\prime}(I^{\prime}):=K(I^{\prime})+\widetilde{K}(I^{\prime}):=K(I^{\prime})+\langle P(I^{\prime},\cdot)\rangle$, and $G=(\partial_{I^{\prime}}K^{\prime})^{-1}\circ K_{I}$. Additionally, setting $(\partial_{I^{\prime}}^{2}K^{\prime}(I^{\prime}))^{-1}:=T(I^{\prime})+\widetilde{T}(I^{\prime})$ for $I^{\prime}\in\mathscr{D}^{\prime}$, the following estimates hold:
$\|\partial_{I^{\prime}}^{2}\widetilde{K}\|_{\overline{r}/2,\mathscr{D}^{\prime}}\leq\mathrm{K}\mathrm{L},\qquad\|G-\text{id}\|_{\overline{r},\mathscr{D}}\leq\overline{r}\mathrm{L},\qquad\|\widetilde{T}\|_{\mathscr{D}^{\prime}}\leq\mathrm{T}\mathrm{L},$ (7)
$\max\left\{\|M\tilde{\phi}\|_{\overline{r}/2,s^{\prime},\mathscr{D}^{\prime}},\|\pi_{2}\partial_{\theta^{\prime}}\tilde{\phi}\|_{\overline{r}/2,s^{\prime},\mathscr{D}^{\prime}}\right\}\leq C_{1}\frac{\varepsilon}{\alpha r\sigma^{\nu}},\quad\|P^{\prime}\|_{\overline{r}/2,s^{\prime},\mathscr{D}^{\prime}}\leq C_{1}\rho\varepsilon,$
with $M:=\text{diag}(r^{-1}\mathbb{I}_{d},\sigma^{-1}\mathbb{I}_{d})$.
For the approximation of smooth functions by real-analytic functions and the uniform convergence of sequences of real-analytic functions, we refer to Lemma 3 and Lemma 4 in [39], respectively.
Theorem 2.17 (Jackson, Moser, Zehnder).
Let $l>0$. There exists a constant $C=C(d,l)>0$ such that for any $f\in C^{l}(\mathbb{R}^{d}\times\mathbb{T}^{d})$ and $s>0$, there is a real-analytic function $f_{s}:\Omega_{s}\to\mathbb{C}$ defined on the complex domain
$\Omega_{s}:=\left\{(I,\theta)\in\mathbb{C}^{d}\times\mathbb{C}^{d}\ \big{|}\ \max\{|\textup{Im}\,I|,|\textup{Im}\,\theta|\}<s\right\},$
satisfying the following:
  1. 1.
    Uniform bound:
    $\sup_{\Omega_{s}}|f_{s}|\leq C\|f\|_{C^{0}}.$
  2. 2.
    Approximation error: For any integer $0\leq l^{\prime}\leq l$,
    $\|f-f_{s}\|_{C^{l^{\prime}}}\leq C\|f\|_{C^{l}}\cdot s^{l-l^{\prime}}.$
  3. 3.
    Derivative stability: For any $0<s^{\prime}<s$ and multi-index $\alpha$ with $|\alpha|\leq l^{\prime}$,
    $\sup_{\Omega_{s^{\prime}}}\left|\partial^{\alpha}f_{s}-\partial^{\alpha}f_{s^{\prime}}\right|\leq C\|f\|_{C^{l}}s^{l-l^{\prime}}.$
Moreover, if $f$ is periodic in a component $I_{i}$ or $\theta_{i}$, then $f_{s}$ preserves periodicity in that component.
Theorem 2.18 (Bernstein, Moser).
Let $\{f_{j}\}_{j\geq 0}$ be a sequence of real-analytic functions defined on nested domains
$\Omega_{j}:=\left\{(I,\theta)\in\mathbb{C}^{d}\times\mathbb{C}^{d}\ \big{|}\ |\textup{Im}(I,\theta)|<s_{j}\right\},$
where $s_{j}=s_{0}\kappa^{j}$ for $s_{0}>0$, $0<\kappa<1$, and $l\in\mathbb{R}^{+}\setminus\mathbb{Z}$. Suppose the sequence satisfies
$\sup_{\Omega_{j}}|f_{j}-f_{j-1}|\leq\Gamma\cdot s_{j-1}^{l}\quad\text{for all }j\geq 1,$
where $\Gamma>0$. Then:
  1. 1.
    Uniform convergence: $f_{j}$ converges uniformly on $\mathbb{R}^{d}\times\mathbb{T}^{d}$ to a limit $f\in C^{l}(\mathbb{R}^{d}\times\mathbb{T}^{d})$.
  2. 2.
    Periodicity preservation: If all $f_{j}$ are periodic in a component $I_{i}$ or $\theta_{i}$, the limit $f$ inherits periodicity in that component.
2.5 Extension of Theorems
In this section, we provide three important lemmas along with their proofs, which serve as the key technical tools for the subsequent analysis. It is important to emphasize that the Hamiltonian systems considered in this paper involve multiplicative noise. This represents a significant departure from the classical setup, where the diffusion term is typically assumed to be constant. Although we extended this to time-varying coefficient noise in [48], where the stochastic term takes the form $\int_{0}^{t}g(s)\,dW_{s}$, representing a linear transformation of Brownian motion, the corresponding solution remains a Gaussian process. In contrast, the setup considered here involves a multiplicative diffusion term, whose diffusion coefficient depends nonlinearly on the current state. This results in the non-Gaussian nature of the solution process.
This non-Gaussian property introduces significant analytical challenges and renders many classical tools from Onsager-Machlup functional theory inapplicable. Specifically, foundational results such as Theorems 2.12 and 2.14 no longer hold in this more general setting.
To overcome these difficulties, we first establish a small ball probability estimate Lemma 2.19, which is a crucial component of our analysis. Using this result, we extend Theorem 2.12 via the conditional expectation comparison principle, leading to Lemma 2.20, tailored to our current setup. Building on this and referencing the results of Gilles [21], we further derive an extended version of Theorem 2.14, formulated as Lemma 2.21. These generalizations form the core technical tools for the subsequent analysis.
Lemma 2.19.
[…]
$\lim\limits_{\epsilon\to 0}\mathbb{P}\left(\|\int_{0}^{1}{g(s,X_{s})}\,{\rm d}W_{s}\|\leq\epsilon\right)\geq e^{-C\epsilon^{-2}},$
where $C>0$ is a constant depending on the dimension.
Proof.
Although the integrand $g(t,X_{t})$ is stochastic and the stochastic integral $\int_{0}^{1}g(s,X_{s})\,\mathrm{d}W_{s}$ is, in general, a nonlinear functional of the Brownian motion and hence not Gaussian distributed, a lower bound for its small-ball probability can still be established under appropriate conditions. Specifically, if $g(t,X_{t})$ satisfies the uniform ellipticity condition and is twice continuously differentiable with bounded derivatives, then the desired estimate follows.
We consider the stochastic process defined by
$A(t):=\int_{0}^{t}g(s,X_{s})\,\mathrm{d}W_{s},$
where $g(s,X_{s})$ is a predictable process with respect to the natural filtration of the Brownian motion $W_{s}$, and satisfies a uniform ellipticity condition. By Itô’s isometry, we have
$\mathbb{E}\left[|A(t)-A(s)|^{2}\right]=\mathbb{E}\left[\int_{s}^{t}\mathrm{Tr}(gg^{\top})(r,X_{r})\,\mathrm{d}r\right].$
Under the uniform ellipticity assumption, there exist constants $0<\lambda\leq\Lambda$ such that
$\lambda(t-s)\leq\mathbb{E}\left[|A(t)-A(s)|^{2}\right]\leq\Lambda(t-s),$
which implies that the trace of the quadratic variation process
$\langle A\rangle_{t}=\int_{0}^{t}gg^{\top}\,\mathrm{d}r$
has linear bounds in time.
Hence, by the Dubins–Schwarz theorem, there exists an $n$-dimensional standard Brownian motion $B_{t}$ and a predictable orthogonal process $U_{t}\in O(n)$ such that
$A(t)=U_{t}\,B_{\langle A\rangle_{t}}.$
Since $U_{t}$ is orthogonal, the Euclidean norm is preserved, i.e., $|U_{t}v|=|v|$, and therefore the supremum norm satisfies
$\|A\|:=\sup_{t\in[0,1]}|A(t)|=\sup_{t\in[0,1]}|B_{\langle A\rangle_{t}}|.$
Introducing the time change $\tau=\langle A\rangle_{t}$, and noting that $\mathrm{d}\tau/\mathrm{d}t\in[\lambda,\Lambda]$, we obtain the bounds
$\sup_{t\in[0,1]}|B_{\langle A\rangle_{t}}|\leq\sup_{s\in[0,\Lambda]}|B_{s}|.$
This yields the inclusion of events:
$\left\{\sup_{s\in[0,\Lambda]}|B_{s}|\leq\epsilon\right\}\subseteq\left\{\sup_{t\in[0,1]}\left|\int_{0}^{t}g(s,X_{s})\,\mathrm{d}W_{s}\right|\leq\epsilon\right\}\quad\text{a.s.}.$
Applying the classical small ball estimate for Brownian motion (Theorem 2.15), there exists a constant $C>0$, depending on the dimension and the upper ellipticity bound $\Lambda$, such that
$\mathbb{P}\left(\|A\|\leq\epsilon\right)\geq\mathbb{P}\left(\sup_{s\in[0,\Lambda]}|B_{s}|\leq\epsilon\right)\sim\exp\left(-\frac{C}{\epsilon^{2}}\right),\quad\text{as }\epsilon\to 0.$
Lemma 2.20.
Suppose that $X_{t}$ is the solution of the Itô stochastic differential equation (2), and that assumptions (C1) and (C2) hold. Define
$A(t):=\int_{0}^{t}g(s,X_{s})\,\mathrm{d}W_{s}.$
Let $f(t)\in L^{2}([0,1],\mathbb{R}^{n})$ be a deterministic function, and define the Itô integral
$I(f):=\int_{0}^{1}\left\langle f(s),\,\mathrm{d}W_{s}\right\rangle.$
Then for supremum norm, we have
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left(I(f)\right)\,\bigg{|}\,\|A\|\leq\epsilon\right]=1.$
Proof.
Since $f(t)$ is deterministic and belongs to $L^{2}([0,1])$, and $W_{t}$ is a standard Brownian motion, it follows from properties of the Itô integral that
$\mathbb{E}\left[I(f)\right]=\mathbb{E}\left[\int_{0}^{1}f(s)\,\mathrm{d}W_{s}\right]=0.$
Since the exponential function is convex, Jensen’s inequality yields
$\mathbb{E}\left(\exp\left(I(f)\right)\,\middle|\,\|A\|\leq\epsilon\right)\geq\exp\left(\mathbb{E}\left[I(f)\,\middle|\,\|A\|\leq\epsilon\right]\right)=\exp(0)=1.$
Hence we have the lower bound
$\mathbb{E}\left(\exp\left(I(f)\right)\,\middle|\,\|A\|\leq\epsilon\right)\geq 1.$
On the other hand, we employ the method of conditional expectation bounding (cf. the proof of Lemma 2.19) to relate the event $\|A\|\leq\epsilon$ to the corresponding small ball event $\|W\|\leq\epsilon^{\prime}$. Under assumption (C2), for any $t\in[0,1]$, we have
$\lambda\leq\|g(t,X_{t})\|_{HS}\leq\Lambda,$
where $\|\cdot\|_{HS}$ denotes the Hilbert-Schmidt norm of matrices. Thus, we obtain the pathwise estimate
$c_{1}\lambda\|W\|\leq\|A\|\leq c_{2}\Lambda\|W\|\quad\text{a.s.}.$
The expression $\|\cdot\|$ denotes the supremum norm, and $c_{1}$ and $c_{2}$ are two constants. Therefore, the event
$\left\{\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right\}\subset\left\{\|A\|\leq\epsilon\right\}\subset\left\{\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right\}\quad\text{a.s.}.$
In particular, we have the inequality
$\mathbb{P}\left(\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right)\leq\mathbb{P}\left(\|A\|\leq\epsilon\right)\leq\mathbb{P}\left(\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right).$
From the small ball asymptotics in Theorem 2.11, it is known that for any deterministic function $f\in L^{2}([0,1],\mathbb{R}^{n})$, we have
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left(\int_{0}^{1}f(s)\,\mathrm{d}W_{s}\right)\,\middle|\,\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right]=1,$
and
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left(\int_{0}^{1}f(s)\,\mathrm{d}W_{s}\right)\,\middle|\,\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right]=1.$
Hence, by the squeeze theorem, we immediately obtain
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left(I(f)\right)\,\middle|\,\|A\|\leq\epsilon\right]=1.$
In fact, since $g$ satisfies the uniform non-degeneracy condition, the event $\{\|A\|\leq\epsilon\}$ is asymptotically equivalent to the small ball event $\{\|W\|\leq\epsilon^{\prime}\}$ as $\epsilon\to 0$. Moreover, since the functional $\exp(I(f))$ is continuous with respect to the underlying Brownian path and equals 1 at the zero path, its conditional expectation under small ball constraints converges to 1. Therefore, by bounding the target expectation between two quantities that both converge to 1, the desired limit follows. ∎
Remark 2.
Under the same assumptions as in the previous Lemma 2.20, for any deterministic function $f(t,x)\in L^{2}([0,1],\mathbb{R}^{n})$ and any $p>1$, we have
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left(I(f)^{p}\right)\,\bigg{|}\,\|A\|\leq\epsilon\right]=1.$
Lemma 2.21.
Let $A(t):=\int_{0}^{t}g(s,X_{s})\,\mathrm{d}W_{s}$, where the matrix-valued function $g(s,X_{s})$ satisfies assumption (C2). Let $f\in L^{2}([0,1]^{2})$ be a symmetric function such that the associated Hilbert–Schmidt operator $K(f)$ is nuclear. Then for supremum norm, we have
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left\{\mathcal{S}(f)\right\}\,\middle|\,\|A\|\leq\epsilon\right]=e^{-\mathrm{Tr}(f)},$
where $\mathcal{S}(f):=2\int_{0<s<t<1}f(s,t)\,\mathrm{d}W_{s}\,\mathrm{d}W_{t}$ denotes the second-order Skorokhod integral associated with the kernel $f$.
Proof.
We begin by applying the method of conditional expectation bounding (cf. the proof of Lemma 2.19) to relate the conditioning event $\|A\|<\epsilon$ to a small-ball event of the form $\|W\|<\epsilon^{\prime}$. Under assumption (C2), we have for all $t\in[0,1]$,
$\lambda\leq\|g(t,X_{t})\|_{HS}\leq\Lambda,$
where $\|\cdot\|_{HS}$ denotes the Hilbert-Schmidt norm of a matrix. This yields the pathwise estimate
$c_{1}\lambda\|W\|\leq\|A\|\leq c_{2}\Lambda\|W\|\quad\text{a.s.},$
and the following inclusion of events:
$\left\{\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right\}\subset\left\{\|A\|\leq\epsilon\right\}\subset\left\{\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right\}\quad\text{a.s.}.$
This leads to bounds on small-ball probabilities
$\mathbb{P}\left(\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right)\leq\mathbb{P}\left(\|A\|\leq\epsilon\right)\leq\mathbb{P}\left(\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right).$
Recall that the second-order Skorokhod integral satisfies the identity
$\mathcal{S}(f)=I^{2}(f)-\mathrm{Tr}(f),$
where $I^{2}(f):=\int_{0}^{1}\int_{0}^{1}f(s,t)\,\mathrm{d}W_{s}\,\mathrm{d}W_{t}$ and $\mathrm{Tr}(f):=\sum_{i,j}\lambda_{i,j}\langle h_{i},h_{j}\rangle$ is the trace of the nuclear operator $K(f)$.
According to Theorem 8 in Gilles [21], we have the exponential identity
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left\{I^{2}(f)\right\}\,\middle|\,\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right]=1,$
and
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left\{I^{2}(f)\right\}\,\middle|\,\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right]=1.$
Then, by the squeeze theorem, we deduce that
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left\{I^{2}(f)\right\}\,\middle|\,\|A\|\leq\epsilon\right]=1,$
which yields
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left\{\mathcal{S}(f)\right\}\,\middle|\,\|A\|\leq\epsilon\right]=e^{-\mathrm{Tr}(f)}.$
3 Onsager-Machlup functional for stochastic Hamiltonian systems
In this section, we derive the Onsager-Machlup functional for Hamiltonian systems by calculating the probability ratio of path perturbations within a small neighborhood of a reference path. The main tools used in the derivation are Girsanov’s theorem, the conditional expectation comparison principle, and the results from Sections 2.4 and 2.5. Our results hold for any finite interval $[0,T]$. However, for simplicity, we define the interval as $[0,1]$ in the following discussion. It is worth noting that, to the best of our knowledge, this is the first Onsager-Machlup functional for multiplicative noise in Euclidean space.
We begin by stating the conditions on the functions $H(\theta,I)$ and $\sigma(t,\theta,I)$ that will be assumed throughout the proof:
  • (C1)
    The Hamiltonian function $H\in C^{3}_{b}(\mathbb{T}^{n}\times\mathbb{R}^{n},\mathbb{R})$ is three times continuously differentiable with all third-order derivatives bounded. Furthermore, its gradient $\nabla H$ is globally Lipschitz continuous with a Lipschitz constant $L$.
  • (C2)
    The diffusion matrix $\sigma\in C^{2}_{b}([0,1]\times\mathbb{T}^{n}\times\mathbb{R}^{n},\mathbb{R}^{n\times n})$ is symmetric, positive definite, and bounded for all $t\in[0,1]$. Moreover, it is Lipschitz continuous with respect to the second and third variables, with Lipschitz constant $L$. In addition, $\sigma$ satisfies a uniform ellipticity condition: there exist constants $0<\lambda\leq\Lambda$ such that
    $\lambda\leq\|\sigma\|_{HS}\leq\Lambda,\quad\forall t\in[0,1],$
    where $\|\cdot\|_{HS}$ denotes the Hilbert–Schmidt norm. Consequently, the inverse $\sigma^{-1}$ exists and is continuous, bounded, and symmetric for all $t\in[0,1]$.
Theorem 3.1.
[…]
$OM(\varphi,\dot{\varphi})$ $=\int_{0}^{1}{\left|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s-\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s,$ (8)
where $\dot{\varphi}(t):=\left(\dot{\varphi_{\theta}}(t),\dot{\varphi_{I}}(t)\right)$, and $|\cdot|$ denotes the Euclidean norm in $\mathbb{R}^{n}$.
Proof.
Let the reference path be given by $\varphi(t)=(\varphi_{\theta}(t),\varphi_{I}(t))$, where $\varphi(t)$ is a definite continuous path, and $\varphi(t)-X(0)\in\mathbb{H}^{1}$. We define the stochastic perturbed solution, denoted as $y(t):=(y_{\theta}(t),y_{I}(t))$, as follows:
$y(t)=\varphi(t)+\int_{0}^{t}\sigma(s,y_{s})\,{\rm d}W_{s}.$ (9)
To simplify the notation in the proof, we introduce the term $W^{\sigma}(t)$, which represents the stochastic perturbation in the system:
$W^{\sigma}(t):=\int_{0}^{t}\sigma(s,y_{s})\,{\rm d}W_{s}.$
We define $\tilde{W}(t)$ as follows. Since $\sigma$ is a predictable process, it can be shown, using Girsanov’s Theorem [24], that under the new probability measure $\tilde{\mathbb{P}}$, the process $\tilde{W}(t)$ is a standard $2n$-dimensional Brownian motion.
$\tilde{W}(t)$ $=W(t)-\int_{0}^{t}\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\,{\rm d}s.$ (10)
[…]
$\mathrm{d}y(t)=J\nabla H(y_{t})\,\mathrm{d}t+\sigma(t,y_{t})\,\mathrm{d}\tilde{W}_{t},$ (11)
[…]
To apply Girsanov’s Theorem and achieve the transformation between the two measures, we define the Radon-Nikodym derivative $\mathcal{R}:=\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}$, which represents the change of measure from $\mathbb{P}$ to $\tilde{\mathbb{P}}$. This derivative is given by an exponential martingale associated with the drift terms, which describes the behavior of the Brownian motion under the new measure after the removal of the drift. The Radon-Nikodym derivative is
$\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}=\exp$ $\left(\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right),\,{\rm d}W_{s}\right\rangle-\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s\right),$
We now aim to compute the transition probability of the system’s path remaining close to the reference path $\varphi(t)$. Using Girsanov’s Theorem, this probability can be expressed as
$\quad\frac{\mathbb{P}\left(\left\|X-\varphi\right\|\leq\epsilon\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}=\frac{\tilde{\mathbb{P}}\left(\left\|y-\varphi\right\|\leq\epsilon\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}=\frac{\tilde{\mathbb{P}}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}$ (12)
$=\frac{\mathbb{E}\left(\mathcal{R}\mathbb{I}_{\left\|W^{\sigma}\right\|\leq\epsilon}\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}=\mathbb{E}\left(\mathcal{R}\big{|}\left\|W^{\sigma}\right\|\leq\epsilon\right)$
$=\exp\left\{-\frac{1}{2}\left(\int_{0}^{1}\left|\sigma^{-1}(t,\varphi_{s})\left(\dot{\varphi_{s}}-J\nabla H(\varphi_{s})\right)\right|^{2}\,{\rm d}s\right)\right\}\times\mathbb{E}\left(\exp\left\{\sum_{i=1}^{4}B_{i}\right\}\bigg{|}\left\|W^{\sigma}\right\|\leq\epsilon\right),$
where $B_{i}$ represents the deviations in the path arising from drift and disturbances, it exhibits stochastic properties. This is further clarified by the following detailed expression:
$B_{1}$ $=\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})J\nabla H(y_{s}),\,{\rm d}W_{s}\right\rangle,$
$B_{2}$ $=-\int_{0}^{1}\left\langle\sigma_{\theta}^{-1}(s,y_{s})\dot{\varphi}_{s},\,{\rm d}W_{s}\right\rangle,$
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s,$
$B_{4}$ $=\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s.$
For the first term $B_{1}$, expand $F(t,y):=\sigma^{-1}(t,y_{t})J\nabla H(y_{t})$ in a Taylor series around $x=\varphi_{t}$ with respect to $W^{\sigma}_{t}:=W^{\sigma}(t)$. The expansion is given by
$F(t,y_{t})=F(t,\varphi)+D_{y}F(t,\varphi)\cdot\left(y-\varphi\right)+O(\|y-\varphi\|^{2}).$
where $D_{y}F\in\mathbb{R}^{2n\times 2n}$ is the Jacobian matrix of $F$ with respect to $y$, and $O(\|y-\varphi\|^{2})$ denotes the higher-order remainder term. The specific expansion for $B_{1}$ can be written as
$B_{1}$ $=\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})J\nabla H(y_{s}),\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle{\left(\sigma^{-1}(s,\varphi_{s})J\nabla H(\varphi_{s})+D_{y}\left(\sigma^{-1}(s,\varphi_{s})J\nabla H(\varphi_{s})\right)W^{\sigma}_{s}+R(s)\right)},\,{\rm d}W_{s}\right\rangle$
$=B_{11}+B_{12}+B_{13},$
where:
$\mathbb{E}\left|R(t)\right|^{2}=\mathcal{O}\left(\mathbb{E}\|W^{\sigma}_{t}\|^{4}\right)=\mathcal{O}(\epsilon^{4}),\quad\text{for }\|A\|\leq\epsilon.$
For the first term $B_{2}$, we proceed with a similar Taylor expansion for $B_{1}$
$B_{2}$ $=-\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})\dot{\varphi}_{s},\,{\rm d}W_{s}\right\rangle$
$=-\int_{0}^{1}\left\langle{\left(\sigma^{-1}(s,\varphi_{s})\dot{\varphi}_{s}+D_{y}\left(\sigma^{-1}(s,\varphi_{s})\dot{\varphi}_{s}\right)W^{\sigma}_{s}+R(s)\right)},\,{\rm d}W_{s}\right\rangle$
$=B_{21}+B_{22}+B_{23}.$
Since $B_{11}$ and $B_{21}$, $B_{12}$ and $B_{22}$, as well as $B_{13}$ and $B_{23}$ exhibit similar structural forms, we analyze each pair jointly in the subsequent discussion. Based on conditions $(C_{1})$, $(C_{2})$, and $\varphi(t)-X(0)\in\mathbb{H}^{1}$, we can demonstrate that the function
$\sigma^{-1}(t,\varphi_{t})J\nabla H(\varphi_{t})~\text{and}~\sigma^{-1}(t,\varphi_{t})\dot{\varphi}_{t}\in L^{2}([0,1],\mathbb{R}^{2n}).$
Therefore, according to Lemma 2.20, we have
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{11}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)=1$ (13)
for all $c\in\mathbb{R}$. And
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{21}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)=1$ (14)
for all $c\in\mathbb{R}$.
Secondly, by applying the matrix version of the chain rule for differentiation, we obtain
$B_{12}$ $=\int_{0}^{1}\left\langle D_{y}\left(\sigma^{-1}(s,\varphi_{s})J\nabla H(\varphi_{s})\right)W^{\sigma}_{s},\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle D_{y}\left(\sigma^{-1}(s,\varphi_{s})\right)J\nabla H(\varphi_{s})W^{\sigma}_{s}+\sigma^{-1}(s,\varphi_{s})JD_{y}\nabla H(\varphi_{s})W^{\sigma}_{s},\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle-\sigma^{-1}(s,\varphi_{s})\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})J\nabla H(\varphi_{s})W^{\sigma}_{s}\right$.
$\quad\left.+\sigma^{-1}(s,\varphi_{s})JD_{y}\nabla H(\varphi_{s})W^{\sigma}_{s},\,{\rm d}W_{s}\right\rangle,$
where $D_{y}\sigma(t,\varphi_{t})$ refers to the derivative tensor of the matrix-valued function $\sigma(t,\varphi_{t})$ with respect to $y_{t}\in\mathbb{R}^{n}$. Specifically, $D_{y}\sigma(t,\varphi_{t})$ is a third-order tensor of dimension $n\times n\times n$, whose components are given by
$(D_{y}\sigma(t,\varphi_{t}))_{ijk}:=\frac{\partial\sigma_{ij}(t,y_{t})}{\partial y_{k}}\bigg{|}_{(t,\varphi_{t})}\quad\text{for }i,j,k=1,\dots,n,$
where $\sigma_{ij}(t,y_{t})$ denotes the $(i,j)$-th element of the matrix $\sigma(t,y_{t})$, and $\frac{\partial\sigma_{ij}(t,y_{t})}{\partial y_{k}}$ is the partial derivative of $\sigma_{ij}(t,y_{t})$ with respect to the $k$-th component of $y_{t}$ (i.e., the different coordinates of $y_{k}$). And
$D_{y}\nabla H(\varphi_{t})=\begin{pmatrix}\frac{\partial^{2}H}{\partial\varphi_{\theta}^{2}}&\frac{\partial^{2}H}{\partial\varphi_{\theta}\partial\varphi_{I}}\\ \frac{\partial^{2}H}{\partial\varphi_{I}\partial\varphi_{\theta}}&\frac{\partial^{2}H}{\partial\varphi_{I}^{2}}\end{pmatrix}$
represents the Hessian matrix of $H$. And
$B_{22}$ $=-\int_{0}^{1}\left\langle D_{y}\left(\sigma^{-1}(s,\varphi_{s})\dot{\varphi}_{s}\right)W^{\sigma}_{s},\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle\sigma^{-1}(s,\varphi_{s})\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\dot{\varphi}_{s}W^{\sigma}_{s},\,{\rm d}W_{s}\right\rangle.$
To simplify the notation, define $F(t,\varphi_{t})\in\mathbb{R}^{n\times n}$ as
$F(t,\varphi_{t})=\sigma^{-1}(t,\varphi_{t})\left[JD_{y}\nabla H(\varphi_{t})-\left(D_{y}\sigma(t,\varphi_{t})\right)\sigma^{-1}(t,\varphi_{t})\left(J\nabla H(\varphi_{t})-\dot{\varphi}_{t}\right)\right].$
Hence, by substituting the expression for $W^{\sigma}(t)$, we obtain
$B_{12}+B_{22}$ $=\int_{0}^{1}\left\langle{F(s,\varphi_{s})W^{\sigma}_{s}},\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle{F(s,\varphi_{s})\int_{0}^{s}{\sigma\left(t,y_{t}\right)}\,{\rm d}W_{t}},\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle{F(s,\varphi_{s})\int_{0}^{s}{\sigma\left(t,\varphi_{t}\right)}\,{\rm d}W_{t}},\,{\rm d}W_{s}\right\rangle$
$\quad+\int_{0}^{1}\left\langle{F(s,\varphi_{s})\int_{0}^{s}\left({\sigma\left(t,y_{t}\right)-\sigma\left(t,\varphi_{t}\right)}\right)\,{\rm d}W_{t}},\,{\rm d}W_{s}\right\rangle$
$=B_{\cdot 21}+B_{\cdot 22}.$
We will examine the term $B_{\cdot 21}$, which can be expressed as a double stochastic integral with respect to $W$:
$B_{\cdot 21}$ $=\int_{0}^{1}\int_{0}^{s}\left\langle{F(s,\varphi_{s})\sigma\left(t,\varphi_{t}\right)}\,{\rm d}W_{t},\,{\rm d}W_{s}\right\rangle.$
Define the function $\bar{F}(s,t)$ as follows:
$\bar{F}(s,t):=F(s,\varphi_{s})\sigma(t,\varphi_{t}).$
Additionally, define the symmetrized matrix of $\bar{F}$, denoted by $\tilde{F}$, as
$\tilde{F}:=\frac{1}{2}(\bar{F}+\bar{F}^{*}).$
Thus, we have
$B_{\cdot 21}=\int_{0}^{1}\int_{0}^{1}\left\langle\tilde{F}\,{\rm d}W_{t},\,{\rm d}W_{s}\right\rangle.$
According to condition $(C2)$, the integral operator K(F~)K($\tilde${F}) associated with the symmetrized kernel $\tilde{F}$ is nuclear, and its trace can be computed as follows:
[…]
$=\int_{0}^{1}\mathrm{Tr}\big{(}\sigma^{-1}(s,\varphi_{s})\left[JD_{y}\nabla H(\varphi_{s})-\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\times\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right]\sigma(s,\varphi_{s})\big{)}\,{\rm d}s$
$=\int_{0}^{1}\mathrm{Tr}\left(JD_{y}\nabla H(\varphi_{s})-\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s.$
Here, we used the cyclic property of the matrix trace: for any three matrices $A$, $B$, and $C$ of compatible dimensions, the identity $\mathrm{Tr}(ABC)=\mathrm{Tr}(BCA)=\mathrm{Tr}(CAB)$ holds. In particular, this allows us to cyclically rearrange the terms inside the trace when computing $\mathrm{Tr}\left(\bar{F}(s,s)\right)$. Additionally, due to the linearity of the trace, $\mathrm{Tr}(A+B)=\mathrm{Tr}(A)+\mathrm{Tr}(B)$, we can analyze $\mathrm{Tr}\left(JD_{y}\nabla H(\varphi_{s})\right)$ separately. Since $J$ is the standard symplectic matrix and $D_{y}\nabla H(\varphi_{s})$ represents the Hessian matrix of $H$, we can prove that $\mathrm{Tr}\left(JD_{y}\nabla H(\varphi_{s})\right)=0$.
By Lemma 2.21, we have
$\quad\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{\cdot 21}\right\}\big{|}\|W^{\sigma}\|<\epsilon\right)$ (15)
$={\rm exp}\left\{\frac{1}{2}\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s\right\}$
for all $c\in\mathbb{R}$. We now analyze the term $B_{\cdot 22}$. Under the condition $\|W^{\sigma}\|\leq\epsilon$, we assume without loss of generality that $\mathbb{P}(\|W^{\sigma}\|\leq\epsilon)=\mathbb{P}(\omega_{1})$, meaning that within the probability event $\omega_{1}$, the condition $\|W^{\sigma}\|\leq\epsilon$ holds. Since $\sigma$ satisfies condition $(C2)$, we can conclude that there exists a constant $C$ such that $\|W\|\leq C\epsilon$ holds within the probability event $\omega_{1}$.
We now consider the integrand in $B_{\cdot 22}$, the term $\int_{0}^{s}\left(\sigma(t,y_{t})-\sigma(t,\varphi_{t})\right)\,\mathrm{d}W_{t}$. Based on the Lipschitz continuity of $\sigma$, we have $\sigma(t,y_{t})-\sigma(t,\varphi_{t})\leq L|y_{t}-\varphi_{t}|=L\|W^{\sigma}\|$. Therefore, we can infer that this term is of order $\mathcal{O}(\epsilon^{2})$ within the event $\omega_{1}$. That is, we can write $B_{\cdot 22}$ as
$B_{\cdot 22}=\int_{0}^{1}\left\langle R(s),\,\mathrm{d}W_{s}\right\rangle,$
where $R(s)$ denotes the remainder terms satisfying $\mathbb{E}|R(s)|^{2}=\mathcal{O}(\epsilon^{4})$. Therefore, we will consider $B_{\cdot 22}$ together with $B_{13}$. Last,
$B_{\cdot 22}+B_{13}+B_{23}$ $=\int_{0}^{1}\left\langle{R(s)},\,{\rm d}W_{s}\right\rangle.$
For any $c_{1},c_{2},c_{3}\in\mathbb{R}$, $c=3\max\left\{c_{1},c_{2},c_{3}\right\}$ and $\delta>0$, we have
$\quad\mathbb{E}\left({\rm exp}\left\{c_{1}B_{122}+c_{2}B_{13}+c_{3}B_{23}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)$ (16)
$\leq\int_{0}^{\infty}{e^{x}\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle{R(s)},\,{\rm d}W_{s}\right\rangle\right|>x\big{|}\|W^{\sigma}\|\leq\epsilon\right)}\,{\rm d}x$
$\leq\int_{\delta}^{\infty}{e^{x}\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle{R(s)},\,{\rm d}W_{s}\right\rangle\right|>x\big{|}\|W^{\sigma}\|\leq\epsilon\right)}\,{\rm d}x+e^{\delta}.$
Define the martingale $M(t)=c\int_{0}^{t}\left\langle{R(s)},\,{\rm d}W_{s}\right\rangle$. We have estimate about its quadratic variation
$\langle M(t)\rangle=c^{2}\int_{0}^{t}{\mathbb{E}\left|R(s)\right|^{2}}\,{\rm d}s\leq C\epsilon^{4}$
for some $C>0$. Using the exponential inequality for martingales, we obtain
$\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle{R(s)},\,{\rm d}W_{s}\right\rangle\right|>x,\|W^{\sigma}\|\leq\epsilon\right)\leq{\rm exp}\left\{-\frac{x^{2}}{2C\epsilon^{4}}\right\}.$
Then, by Lemma 2.19, we have
$\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle{R(s)},\,{\rm d}W_{t}\right\rangle\right|>x\big{|}\|W^{\sigma}\|\leq\epsilon\right)\leq{\rm exp}\left\{-\frac{x^{2}}{2C\epsilon^{4}}\right\}{\rm exp}\left\{\frac{C}{\epsilon^{2}}\right\}.$
[…]
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{c_{1}\bar{B}_{22}+c_{2}B_{13}+c_{3}B_{23}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)=1$ (17)
for all $c_{1},c_{2},c_{3}\in\mathbb{R}$, as $\epsilon\to 0$ and $\delta\to 0$.
For the third term $B_{3}$, let $v_{t}:=J\nabla H(y_{t})-\dot{\varphi}_{t}$, then
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left(\left|\sigma^{-1}(s,\varphi_{s})v_{s}\right|^{2}-\left|\sigma^{-1}(s,y_{s})v_{s}\right|^{2}\right)\,{\rm d}s$
$=\frac{1}{2}\int_{0}^{1}\left\langle\left(\left({\sigma^{-1}\left(s,y_{s}\right)}\right)^{2}-\left({\sigma^{-1}\left(s,\varphi_{s}\right)}\right)^{2}\right)v_{s},v_{s}\right\rangle\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left(\left\|\left({\sigma^{-1}\left(s,y_{s}\right)}\right)^{2}-\left({\sigma^{-1}\left(s,\varphi_{s}\right)}\right)^{2}\right\|_{2}\right)\left|v_{s}\right|^{2}}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left\|{\sigma^{-1}\left(s,y_{s}\right)}-{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left\|{\sigma^{-1}\left(s,y_{s}\right)}+{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left|v_{s}\right|^{2}}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}\left\|{\sigma^{-1}\left(s,y_{s}\right)}\right\|_{2}\left\|{\sigma\left(s,\varphi_{s}\right)-\sigma\left(s,y_{s}\right)}\right\|_{2}\left\|{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}$
$\qquad\times\left\|{\sigma^{-1}\left(s,y_{s}\right)}+{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left|v_{s}\right|^{2}\,{\rm d}s.$
Since both $\sigma^{-1}$ and $v$ are bounded, and $\sigma$ is Lipschitz continuous, while $\varphi_{t}-X_{0}\in\mathbb{H}^{1}$, we conclude that there exists a constant $C>0$ such that
$B_{3}\leq C\int_{0}^{1}\left\|\sigma(s,\varphi_{s})-\sigma(s,y_{s})\right\|_{2}\,{\rm d}s\leq CL\left\|W^{\sigma}\right\|.$
The above inequality implies that
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{3}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)=1$ (18)
for all $c\in\mathbb{R}$.
For the fourth term $B_{4}$, let $v_{t}:=J\nabla H(y_{t})-\dot{\varphi}_{t}$, $u:=J\nabla H(\varphi_{t})-\dot{\varphi}_{t}$, then
$B_{4}$ $=\frac{1}{2}\int_{0}^{1}{\left|\sigma^{-1}(s,\varphi_{s})u\right|^{2}-\left|\sigma^{-1}(s,\varphi_{s})v\right|^{2}}\,{\rm d}s$
$=\frac{1}{2}\int_{0}^{1}{\left\langle\sigma^{-1}(s,\varphi_{s})\left(u-v\right),\sigma^{-1}(s,\varphi_{s})\left(u+v\right)\right\rangle}\,{\rm d}s$
$=\frac{1}{2}\int_{0}^{1}{\left\langle\sigma^{-1}(s,\varphi_{s})^{2}\left(u-v\right),\left(u+v\right)\right\rangle}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left\|\sigma^{-1}(s,\varphi_{s})^{2}\right\|_{2}\left\|\left(u-v\right)\right\|\left\|\left(u+v\right)\right\|}\,{\rm d}s.$
Since both $\sigma^{-1}$, $v$, and $u$ are bounded, and $J\nabla H(y_{t})$ is Lipschitz continuous, and noting that $\varphi_{t}-x_{0}\in\mathbb{H}^{1}$, we can conclude that there exists a constant $C>0$ such that
$B_{4}\leq C\left|u-v\right|\leq C\int_{0}^{1}\left|J\nabla H(\varphi_{s})-J\nabla H(y_{s})\right|\,{\rm d}s\leq CL\left\|W^{\sigma}\right\|.$
So we have
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{4}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)=1$ (19)
for all $c\in\mathbb{R}$.
[…]
$\lim\limits_{\epsilon\to 0}\frac{\mathbb{P}\left(\|X_{t}-\varphi_{t}\|\leq\epsilon\right)}{\mathbb{P}\left(\|W\|\leq\epsilon\right)}$ $={\rm exp}\left\{-\frac{1}{2}\int_{0}^{1}{\left|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s\right$.
$\qquad\quad\left.+\frac{1}{2}\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s\right\},$
where the term $D_{y}\sigma(t,\varphi_{t})$ refers to the third-order $n\times n\times n$ derivative tensor of the matrix-valued function $\sigma(t,y_{t})$ with respect to $y_{t}\in\mathbb{R}^{n}$, evaluated at $\varphi_{t}$. Consequently,
$OM(\varphi,\dot{\varphi})=\int_{0}^{1}{\left|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s-\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s.$
4 Most Probable Path
In this section, building upon Theorem 3.1, we further introduce the method of Euler–Lagrange equations to study the most probable paths in the solution trajectories of stochastic differential equations. It is well known that, under small noise limits or small-ball probability constraints, the maximum likelihood path of a stochastic process tends to minimize the so-called Onsager-Machlup functional, given by
$OM(\varphi,\dot{\varphi})=\int_{0}^{1}{\left|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s-\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s.$
The integrand
$\left|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}-\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)$
can be viewed as the Lagrangian function associated with this variational problem. According to the classical calculus of variations, the optimal path $\varphi$ should satisfy the corresponding Euler–Lagrange equation
$\frac{\partial L}{\partial\varphi}-\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\partial L}{\partial\dot{\varphi}}\right)=0,$
together with the initial condition $\varphi_{0}=X_{0}$ and an appropriate terminal condition.
Due to the presence of nonlinear matrix-valued functions, trace operators, and nontrivial drift structures, the above equation is typically challenging to solve explicitly. However, it plays a critical role in both theoretical analysis and numerical investigations. To address this challenge, we adopt the Ansatz method as a more tractable and analytically convenient approach. Accordingly, we present the following theorem on the solution of the maximum probability path.
Theorem 4.1.
[…]
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}.$ (20)
Proof.
Based on the Ansatz method, we first consider a simplified setting in which the term $\left(D_{y}\sigma(t,\varphi_{t})\right)$ is assumed to depend solely on time $t$ and is denoted by $\mathcal{D}^{\sigma}(t)$. Under this simplification, the Onsager-Machlup functional can be reduced to the following form:
$OM(\varphi,\dot{\varphi})=\int_{0}^{1}{\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s-\int_{0}^{1}\mathrm{Tr}\left(\mathcal{D}^{\sigma}(s)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s.$ (21)
To facilitate the optimization, we introduce the auxiliary variable
$A(t):=\sigma^{-1}(t,\varphi_{t})\left(J\nabla H(\varphi_{t})-\dot{\varphi_{t}}\right),$
which reformulates the Onsager-Machlup functional as
$OM(A)=\int_{0}^{1}\left(\|A(s)\|^{2}-\mathrm{Tr}(\mathcal{D}^{\sigma}(s)A(s))\right)ds.$
Since $\mathcal{D}^{\sigma}(t)$ is independent of $\varphi$, this expression corresponds to a variational problem where the integrand depends quadratically and linearly on the auxiliary variable $A(t)$. Given that the Frobenius norm is both differentiable and convex, we can compute the extremum by taking the first variation with respect to $A$.
For each fixed $t$, we define the Lagrangian function
$\mathcal{L}(A):=\|A\|^{2}-\mathrm{Tr}(\mathcal{D}^{\sigma}A),$
where $\mathcal{D}^{\sigma}(t):=\mathcal{D}^{\sigma}(t)_{ijk}$ denotes a time-dependent rank-3 tensor and $A:=A_{j}$ is an $n$-dimensional vector. The first variation of $\mathcal{L}$ in the direction of $\delta A$ is given by
$\delta\mathcal{L}$ $=2A^{\top}\delta A-\mathrm{Tr}(\left(\mathcal{D}^{\sigma}\right)^{\top}\cdot\delta A)$
$=2\sum_{j=1}^{n}A_{j}\delta A_{j}-\sum_{i,j}\mathcal{D}^{\sigma}_{jii}\delta A_{j}$
$=\sum_{j=1}^{n}\left(2A_{j}-\sum_{i}\mathcal{D}^{\sigma}_{jii}\right)\delta A_{j},$
which implies that the unique minimizer is
$A^{*}_{j}=\frac{1}{2}\sum_{i}D_{jii}.$
Substituting back the definition of $A$, we obtain the expressions for the most probable path $\varphi(t)$ in the form of a drift-corrected Hamiltonian system
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\dfrac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left({\mathcal{D}^{\sigma}}(t)\right)_{\cdot ii}.$
Based on the Ansatz method, we now assume that the spatial derivatives of the diffusion coefficients, $D_{y}\sigma(t,\varphi_{t})$ is equal to their evaluations at a reference path $\psi(t)=(\psi_{\theta}(t),\psi_{I}(t))$, namely,
$D_{y}\sigma(t,\varphi_{t})=D_{y}\sigma(t,\psi_{t}).$
We further assume that the reference path $\psi(t)$ satisfies the following drift-corrected Hamiltonian system:
$\dot{\psi}(t)=J\nabla H(\psi_{t})-\frac{1}{2}\sigma(t,\psi_{t})\sum_{i}\left({D_{y}}\sigma(t,\psi_{t})\right)_{\cdot ii}.$
Under this assumption, we may again abbreviate the tensor $D_{y}\sigma(t,\psi_{t})$ by the time-dependent third-order tensors $\mathcal{D}^{\sigma}(t)$, which is treated as independent of the variable $\varphi=\left(\varphi_{I},\varphi_{\theta}\right)$.
Following the same steps as in the earlier derivation, we obtain that the most probable path $\varphi(t)$, which minimizes the Onsager-Machlup functional, satisfies the following system:
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\psi_{t})\right)_{\cdot ii}.$
Since both $\varphi(t)$ and $\psi(t)$ satisfy the same system with identical initial conditions, the uniqueness of solutions to ordinary differential equations implies that $\psi(t)=\varphi(t)$. Therefore, the path $\varphi(t)$ that minimizes the Onsager-Machlup functional - the most probable path - must satisfy the following drift-corrected Hamiltonian system:
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}.$
This system represents a deterministic perturbation of the classical Hamiltonian equations, where the correction terms arise from the effect of multiplicative noise. In particular, the additional term involving $\frac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}$ quantify the influence of the multiplicative noise geometry (represented through the spatial derivative tensors of the diffusion matrices) on the most probable evolution path. Under certain regularity assumptions, such as the invertibility of $\sigma$, as well as the smoothness of the Hamiltonian $H\in C_{b}^{3}$, the solution corresponding to the most probable path $\varphi(t)$ is uniquely determined by the initial conditions and the modified dynamics.
Moreover, we observe that the most probable path of the stochastic Hamiltonian system (20) does not, in general, preserve the Hamiltonian structure. To address this issue, we introduce Condition $(C3)$, which is a slightly stronger version of Condition $(C2)$. In Condition $(C3)$, the symmetric matrix assumed in $(C2)$ is further restricted to be diagonal. This modification is both natural and intuitive: while symmetry is a necessary requirement in the derivation of the Onsager-Machlup functional, the preservation of Hamiltonian structure inherently demands antisymmetry. Furthermore, to ensure the conservation of the Hamiltonian, we assume that the matrix $\sigma$ is independent of time $t$. Hence, Condition $(C3)$ proposes the use of a diagonal matrix $\sigma$, which is simultaneously symmetric and antisymmetric.
  • (C3)
    The diffusion matrix $\sigma(\theta,I)\in C^{2}(\mathbb{R}^{n}\times\mathbb{R}^{n},\mathbb{R}^{n\times n})$ is diagonal, positive definite, and bounded for all $t\in[0,1]$. Additionally, it is Lipschitz continuous with respect to the second and third variables, with Lipschitz constant $L$. Furthermore, $\sigma$ satisfies a uniform ellipticity condition: there exist constants $0<\lambda\leq\Lambda$ such that
    $\lambda\leq\|\sigma\|_{HS}\leq\Lambda,\quad\forall t\in[0,1],$
    where $\|\cdot\|_{HS}$ denotes the Hilbert–Schmidt norm. Consequently, the inverse $\sigma^{-1}$ exists and is continuous, bounded, and symmetric for all $t\in[0,1]$.
Theorem 4.2.
Assuming that Conditions $(C1)$ and $(C3)$ are satisfied, we further demonstrate, based on the established Onsager-Machlup functional, that the maximum probability path of the stochastic Hamiltonian system (2) satisfies the following nearly integrable Hamiltonian system:
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}J\nabla H_{\sigma}(\varphi_{t}),$ (22)
where $H_{\sigma}(\varphi_{t}):=\frac{1}{2}\sum_{i=1}^{n}\left(\sigma_{\theta,i}(\theta,I)^{2}-\sigma_{I,i}(\theta,I)^{2}\right)$, $\sigma_{\theta}(\theta,I)=\left(\sigma_{\theta,i}(\theta,I)\right)_{i=1}^{n}$ and $\sigma_{I}(\theta,I)=\left(\sigma_{I,i}(\theta,I)\right)_{i=1}^{n}$.
Proof.
[…]
$\sigma(\theta,I):=\begin{pmatrix}\sigma_{\theta}(\theta,I)\!\!&\!\!0\\ 0\!\!&\!\!\sigma_{I}(\theta,I)\end{pmatrix}=\begin{pmatrix}\sigma_{\theta,1}&0&\cdots&0&0&0&\cdots&0\\ 0&\sigma_{\theta,2}&\cdots&0&0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\sigma_{\theta,n}&0&0&\cdots&0\\ 0&0&\cdots&0&\sigma_{I,1}&0&\cdots&0\\ 0&0&\cdots&0&0&\sigma_{I,2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0&0&0&\cdots&\sigma_{I,n}\\ \end{pmatrix}$,
and
$\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}=\begin{pmatrix}\frac{\partial\sigma_{\theta,1}}{\partial\theta_{1}}&\frac{\partial\sigma_{\theta,2}}{\partial\theta_{2}}&\cdots&\frac{\partial\sigma_{\theta,n}}{\partial\theta_{n}}&\frac{\partial\sigma_{I,1}}{\partial I_{1}}&\frac{\partial\sigma_{I,2}}{\partial I_{2}}&\cdots&\frac{\partial\sigma_{I,n}}{\partial I_{n}}\end{pmatrix}^{T}.$
Thus,
$\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}=\begin{pmatrix}\sigma_{\theta,1}\frac{\partial\sigma_{\theta,1}}{\partial\theta_{1}}&\sigma_{\theta,2}\frac{\partial\sigma_{\theta,2}}{\partial\theta_{2}}&\cdots&\sigma_{\theta,n}\frac{\partial\sigma_{\theta,n}}{\partial\theta_{n}}&\sigma_{I,1}\frac{\partial\sigma_{I,1}}{\partial I_{1}}&\sigma_{I,2}\frac{\partial\sigma_{I,2}}{\partial I_{2}}&\cdots&\sigma_{I,n}\frac{\partial\sigma_{I,n}}{\partial I_{n}}\end{pmatrix}^{T}.$
Based on the above relations, it is straightforward to verify that the resulting system can be expressed in a Hamiltonian structure. The Hamiltonian can be defined as
$H_{\sigma}(\varphi_{t}):=\frac{1}{2}\sum_{i=1}^{n}\left(\sigma_{\theta,i}(\theta,I)^{2}-\sigma_{I,i}(\theta,I)^{2}\right).$
Remark 3.
[…]
$\sum_{k}\sigma_{ik}\,\frac{\partial}{\partial\varphi_{j}}\Bigg{(}\sum_{m}\frac{\partial\sigma_{km}}{\partial\varphi_{m}}\Bigg{)}=\sum_{k}\sigma_{jk}\,\frac{\partial}{\partial\varphi_{i}}\Bigg{(}\sum_{m}\frac{\partial\sigma_{km}}{\partial\varphi_{m}}\Bigg{)}$
If this condition holds, it can be shown that the term $\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}$ can be formulated in the Hamiltonian structure. However, this stringent condition is intricate and challenging to verify in practice. Thus, it is not adopted in this work. In contrast, Condition $(C3)$ provides a more practical and tractable framework, making it a particularly favorable choice for analysis.
Corollary 4.3.
We consider here a special case, namely a stochastic Hamiltonian system with additive noise, whose dynamics are governed by the following stochastic differential equation:
$\mathrm{d}X(t)=J\nabla H(X(t))\,\mathrm{d}t+\sigma(t)\,\mathrm{d}W(t),$
where the diffusion coefficient $\sigma(t)$ depends only on the time variable $t$ and is independent of the state variable. According to Theorem 3.1, the Onsager-Machlup functional corresponding to this system can be derived as
$OM(\varphi,\dot{\varphi})$ $=\int_{0}^{1}\left|\sigma^{-1}(s)\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s.$
Moreover, by applying the Euler–Lagrange equation, we obtain that the most probable path $X^{*}(t)$ satisfies
$\mathrm{d}{X^{*}}(t)=J\nabla H(X^{*}(t))\,\mathrm{d}t.$
This result shows that, in the case of additive noise, the most probable path coincides exactly with the trajectory of the corresponding deterministic Hamiltonian system, without any additional drift correction induced by the noise. This behavior is in sharp contrast to the case of multiplicative noise, where stochastic perturbations generally introduce nontrivial modifications to the drift. Therefore, for a classical Hamiltonian system, the presence of additive noise does not alter the evolution of its most probable trajectory.
Corollary 4.4.
We consider a completely integrable stochastic Hamiltonian system, which is expressed as follows:
$\begin{cases}\,{\rm d}\theta_{i}(t)=\frac{\partial H_{0}}{\partial I_{i}}(\theta,I)\,{\rm d}t+\sum_{k=1}^{n}\frac{\partial H_{k}}{\partial I_{i}}(\theta,I)\,{\rm d}W_{k}(t),\\ \,{\rm d}I_{i}(t)=-\frac{\partial H_{0}}{\partial\theta_{i}}(\theta,I)\,{\rm d}t-\sum_{k=1}^{n}\frac{\partial H_{k}}{\partial\theta_{i}}(\theta_{i},I)\,{\rm d}W_{k}(t),\end{cases}$ (23)
where $\{H_{k}\}_{k=0}^{n}$ denotes a family of completely integrable Hamiltonian functions. The quantities $\theta_{i}(t)$ and $I_{i}(t)$ are, respectively, the $i$-th components of the angle vector $\theta(t)$ and the action vector $I(t)$, for $1\leq i\leq n$; $W_{k}(t)$ is independent one-dimensional standard Brownian motion for $1\leq k\leq n$. Consequently, system (23) can be written equivalently as
$\mathrm{d}X(t)=J\nabla H_{0}(X(t))\,\mathrm{d}t+J\nabla\bar{H}(X(t))\,\mathrm{d}W(t),$ (24)
where $J$ denotes the standard symplectic matrix, and $\nabla H_{0}(X)$ is the gradient of the Hamiltonian $H_{0}$ with respect to $X=(\theta,I)$. The notation $\nabla\bar{H}$ refers to the $2n\times n$ matrix whose columns are the gradients of the Hamiltonians $H_{1},H_{2},\dots,H_{n}$, namely,
$\nabla\bar{H}:=\left(\nabla H_{1}(X),\,\nabla H_{2}(X),\,\dots,\,\nabla H_{n}(X)\right).$
The process $W(t)$ is a $n$-dimensional standard Brownian motion, and all these Brownian motions are mutually independent. Due to the complete integrability of the system, for all $0\leq k\leq n$, we have $\frac{\partial H_{k}}{\partial\theta}(\theta,I)=0$. It is readily observed that condition $(C2)$ is no longer valid. However, this does not affect our conclusions, as equation (23) can be reformulated as follows:
$\begin{cases}\,{\rm d}\theta_{i}(t)=\frac{\partial H_{0}}{\partial I_{i}}(I)\,{\rm d}t+\sum_{k=1}^{n}\frac{\partial H_{k}}{\partial I_{i}}(I)\,{\rm d}W_{k}(t),\\ \,{\rm d}I_{i}(t)=0.\end{cases}$ (25)
Since the equation $\mathrm{d}I_{i}(t)=0$ is deterministic, we only need to analyze $\theta_{i}(t)$ separately. We may assume that $\frac{\partial H_{k}}{\partial I_{i}}(I)\neq 0$, which is a standard assumption. This, in particular, implies that the matrix $\sigma:=\left(\frac{\partial H_{k}}{\partial I_{i}}(I)\right)_{1\leq i,k\leq n}$ is an invertible $n\times n$ matrix. Therefore, based on Theorem 3.1, we derive the Onsager-Machlup functional for system (25)
$OM(\varphi_{\theta},\dot{\varphi_{\theta}})=\int_{0}^{1}\left\|\sigma^{-1}\left(\frac{\partial H_{0}}{\partial I}(I)-\dot{\varphi_{\theta}}\right)\right\|^{2}\,{\rm d}s.$
Furthermore, by applying the Euler-Lagrange equation, we demonstrate that the most probable path $X^{*}(t)$ satisfies the deterministic Hamiltonian equation
$\mathrm{d}{X^{*}}(t)=J\nabla H_{0}(X^{*}(t))\,\mathrm{d}t.$
This significant finding provides multi-dimensional insights into integrable systems under random perturbations. First, it offers rigorous mathematical verification of Li’s averaging principle [30] from the perspective of probabilistic variational principles, revealing the dominant role of deterministic dynamics in stochastically perturbed systems. Second, it establishes a theoretical connection between the Onsager-Machlup functional and the averaged equations, bridging probabilistic methods with deterministic analysis. Finally, the proposed methodology extends the traditional averaging theory to the realm of variational principles, providing a novel analytical framework for studying more general classes of stochastically perturbed Hamiltonian systems.
Therefore, although stochastic perturbations introduce complexity and uncertainty into Hamiltonian systems, minimizing the Onsager-Machlup functional allows us to identify the dominant factors governing system behavior, namely, the evolution along the most probable path. This finding is of significant importance in both theoretical research and practical applications. To further substantiate this conclusion, we present a specific example: the one-dimensional harmonic oscillator.
Example 4.5.
Consider a one-dimensional harmonic oscillator with the Hamiltonian $H(q,p)=\frac{p^{2}}{2m}+\frac{1}{2}kq^{2}$, where $\frac{p^{2}}{2m}$ represents the kinetic energy term, and $V(q)=\frac{1}{2}kq^{2}$ is the potential energy term. Here, $m=1$ denotes the mass, $k=1$ denotes the spring constant, and $q$ and $p$ respectively represent position and momentum. The classical Hamiltonian equations are given by
$\begin{cases}\begin{aligned} dq(t)&=p\,{\rm d}t,\quad&q(0)&=100,\\ dp(t)&=-q\,{\rm d}t,\quad&p(0)&=100.\end{aligned}\end{cases}$ (26)
Additionally, consider a one-dimensional harmonic oscillator under the influence of white noise perturbations, defined by the following stochastic Hamiltonian system:
$\begin{cases}\begin{aligned} dq(t)&=p\,{\rm d}t+(3+\sin(t)+\cos(\sqrt{2}(q+p)))\,{\rm d}W_{q}(t),&q(0)&=100,\\ dp(t)&=-q\,{\rm d}t+(3+\cos(t)+\sin(\sqrt{3}(q+p)))\,{\rm d}W_{p}(t),&p(0)&=100,\end{aligned}\end{cases}$ (27)
where $W_{q}(t)$ and $W_{p}(t)$ are independent one-dimensional Brownian motions.
The system above clearly satisfies the conditions $(C1)$ and $(C2)$. Therefore, based on Theorem 3.1, we can derive the corresponding Onsager-Machlup functional, and, using the proofs provided earlier in this section, obtain the solution $X^{*}(t)$ for the most probable path of this stochastic Hamiltonian system, which satisfies the following equation:
$\begin{cases}\begin{aligned} dq(t)&=\left(p-2\sqrt{2}\sin(\sqrt{2}(q+p))(3+\sin(t)+\cos(\sqrt{2}(q+p)))\right)\,{\rm d}t,&q(0)&=100,\\ dp(t)&=-\left(q-2\sqrt{3}\cos(\sqrt{3}(q+p))(3+\cos(t)+\sin(\sqrt{3}(q+p)))\right)\,{\rm d}t,&p(0)&=100.\end{aligned}\end{cases}$ (28)
In Figure 1, we clearly observe the distinct contrast between the phase space trajectories of the classical harmonic oscillator, the stochastically perturbed oscillator, and the most probable phase space trajectory. In the absence of stochastic perturbations, the classical harmonic oscillator follows a stable, closed circular orbit, where momentum $p$ and position $q$ oscillate periodically, reflecting energy conservation. This circular trajectory signifies the steady exchange between kinetic and potential energy, thereby maintaining a constant Hamiltonian.
Figure 1: Comparison of phase-space trajectories among the deterministic (black bold solid line), stochastically perturbed (gray thin dashed lines), and most probable (red bold solid line) harmonic oscillators. The system is modeled with the following parameters: mass $m=1$, spring constant $k=1$, time step $dt=0.0001$, and total simulation time $T=500$. The trajectory of the classical system in phase space follows a stable circular orbit, indicating energy conservation and periodic motion. In contrast, the stochastic system incorporates multiplicative noise perturbations through the Euler-Maruyama method, resulting in trajectories that gradually diffuse away from the original orbit, exhibiting both dispersion and irregularity over time. The most probable trajectory, derived from the Onsager-Machlup functional, is located in the densest region of the stochastic trajectories, aligning with the theoretical predictions presented in this study.
However, when multiplicative noise perturbations are introduced, the trajectory deviates from this ideal path, exhibiting diffusion and irregularity. As illustrated in Figure 1, the stochastic oscillator’s trajectory gradually drifts away from the stable orbit, with random fluctuations disrupting the original pattern. This diffusion phenomenon arises because the noise induces energy fluctuations, causing deviations in $p$ and $q$, preventing the system from strictly adhering to the classical oscillator’s trajectory. Although some periodic behavior persists, the perturbations induce instability, leading to a trajectory that expands outward in phase space over time.
Despite these fluctuations, we also observe that the stochastic oscillator’s phase space trajectories remain densely concentrated near the most probable phase space trajectory. Notably, the most probable phase space trajectory remains closely aligned with the deterministic oscillator’s path, albeit with slight deviations caused by the multiplicative noise. Consequently, the most probable trajectory is not strictly conservative. This observation is consistent with our theoretical Theorems 3.1 and Section 4, further validating the conclusion that, even under stochastic perturbations, the system tends to evolve near the deterministic solution, albeit with slight deformations.
5 Large Deviation Principle and Rate Function for Stochastic Hamiltonian Systems
In this section, we derive the large deviation principle for stochastic Hamiltonian systems, combining the Onsager-Machlup functional and Freidlin-Wentzell theory [28]. Large deviation principle describes the probability behavior of a system’s trajectory deviating from the most probable path, with the rate function quantifying the decay rate of this probability.
Since our objective is to study the behavior of the trajectories in the stochastic Hamiltonian system as the noise term approaches zero, we simplify equation (2) into the following form for convenience:
$\mathrm{d}X(t)=J\nabla H(X(t))\,\mathrm{d}t+\epsilon\sigma(t,X(t))\,\mathrm{d}W(t),$ (29)
where $J$ is the standard symplectic matrix, and $\nabla H(X)$ denotes the gradient of the Hamiltonian with respect to $X=(\theta,I)$. The matrix $\sigma(t,X)$ is the $2n\times 2n$ diffusion matrix, and $W(t)$ is the $2n$-dimensional standard Brownian motion. The parameter $\epsilon$ represents the noise intensity. Our primary focus is on analyzing the statistical behavior of this system as $\epsilon\to 0$.
Theorem 5.1.
Assuming that Conditions $(C1)$ and $(C2)$ hold, let $X^{\epsilon}(t):=(\theta(t),I(t))$ be the solution to the following stochastic Hamiltonian system (29). As $\epsilon\to 0$, the most probable path $\varphi^{*}(t)=(\varphi^{*}_{\theta}(t),\varphi^{*}_{I}(t))$ is given by the deterministic Hamiltonian equations
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}\epsilon^{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}.$
For any path $X^{\epsilon}(t)$, the probability that the system deviates from the most probable path satisfies the large deviation principle
$\epsilon^{2}\ln\mathbb{P}(X^{\epsilon}(t)\in A)\approx-\inf_{\varphi\in A}S(\varphi),$ (30)
where $\varphi\in A$ denotes an arbitrarily continuous function, and $A\subset\mathbb{R}^{d}$ denotes an arbitrary measurable set, and suppose that $\varphi^{*}(t)\notin A$. Furthermore, the rate function $S(\varphi)$ is given by
[…]
with $\sigma^{-1}(t,\varphi_{t})$ being the inverses of the diffusion matrices $\sigma(t,\varphi_{t})$.
Proof.
[…]
$\mathrm{d}y(t)=\varphi(t)+\epsilon\int_{0}^{t}\sigma(s,y_{s})\,\mathrm{d}W(s).$
To simplify the notation in the proof, we introduce the term $W^{\sigma}(t)$, which encapsulates the stochastic perturbation in the system
$W^{\sigma}(t):=\int_{0}^{t}\sigma(s,y_{s})\,\mathrm{d}W(s).$
Then we introduce a new probability measure $\tilde{\mathbb{P}}$, under which the transformed Brownian motions are given by
$\tilde{W}(t)$ $=W(t)-\frac{1}{\epsilon}\int_{0}^{t}\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi_{s}}\right)\,{\rm d}s.$
Under this new measure $\tilde{\mathbb{P}}$, the system is purely driven by the new Brownian motion $\tilde{W}(t)$, free of the deterministic drift terms.
The Radon-Nikodym derivative $\mathcal{R}:=\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}$, representing the change of measure from $\mathbb{P}$ to $\tilde{\mathbb{P}}$, is given by the exponential martingale associated with the removed drift terms. This can be written in the following form:
$\mathcal{R}=\exp\left\{\frac{1}{\epsilon}\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi_{s}}\right),\,{\rm d}W(s)\right\rangle-\frac{1}{2\epsilon^{2}}\int_{0}^{1}\left|\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi_{s}}\right)\right|^{2}\,{\rm d}s\right\}.$
Define
$K(\varphi,\epsilon)=\{x-x_{0}\in\mathbb{H}^{1}\mid\varphi-x_{0}\in\mathbb{H}^{1},\|x-\varphi\|\leq\epsilon,\epsilon>0\}.$
Utilizing Girsanov’s Theorem, we aim to compute the probability that the trajectory of the solution $X(t)$ of the stochastic Hamiltonian system remains in close proximity to a reference path $\varphi(t)$ when the noise intensity $\epsilon$ is minimal. Then
$\quad\mathbb{P}(X(t)\in K(\varphi,\epsilon))=\mathbb{P}\left(\left\|X(t)-\varphi(t)\right\|\leq\epsilon\right)=\tilde{\mathbb{P}}\left(\left\|y(t)-\varphi(t)\right\|\leq\epsilon\right)$ (32)
$=\mathbb{E}\left(\mathcal{R}\mathbb{I}_{\left\|W^{\sigma}\right\|\leq 1}\right)=\mathbb{E}\left(\mathcal{R}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)\times\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1)$
$=\exp\left\{-\frac{1}{2\epsilon^{2}}\left(\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}\,{\rm d}t\right)\right\}\times\mathbb{E}\left(\exp\left\{\frac{1}{\epsilon^{2}}\sum_{i=1}^{4}B_{i}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)\times\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1),$
where $B_{i}$ represents the deviations in the path arising from drift and disturbances, it exhibits stochastic properties. This is elaborated upon in the following detailed expressions:
$B_{1}$ $=\epsilon\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})J\nabla H(y_{s}),\,{\rm d}W_{s}\right\rangle,$
$B_{2}$ $=-\epsilon\int_{0}^{1}\left\langle\sigma_{\theta}^{-1}(s,y_{s})\dot{\varphi}_{s},\,{\rm d}W_{s}\right\rangle,$
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s,$
$B_{4}$ $=\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s.$
We decompose the system’s probability into a deterministic part
$\exp\left\{-\frac{1}{2\epsilon^{2}}\left(\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}\,{\rm d}s\right)\right\}=\exp\left\{-\frac{1}{2\epsilon^{2}}\left(\int_{0}^{1}U(\varphi)\,{\rm d}s\right)\right\}$
a correction part
[…]
and small ball probabilities for Brownian motion
$\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1).$
Specifically, the deterministic component represents the behavior along the deterministic trajectory, while the correction component accounts for deviations induced by random perturbations. Large deviation theory focuses on large-scale deviations, and in this context, we are particularly concerned with the $\frac{1}{\epsilon^{2}}$ scale. Given the boundedness of $\sigma^{-1}(s,y)$, along with the fact that $H\in C^{3}_{b}(\mathbb{R}^{n}\times\mathbb{R}^{n},\mathbb{R})$, we can infer that
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{1}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)=1,$ (33)
and
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{2}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)=1.$ (34)
For the third term $B_{3}$, let $v:=J\nabla H(y_{s})-\dot{\varphi}_{s}$, then
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left(\left|\sigma^{-1}(s,\varphi_{s})v\right|^{2}-\left|\sigma^{-1}(s,y_{s})v\right|^{2}\right)\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left(\left\|\left({\sigma^{-1}\left(s,y_{s}\right)}\right)^{2}-\left({\sigma^{-1}\left(s,\varphi_{s}\right)}\right)^{2}\right\|_{HS}\right)\left|v\right|^{2}}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left\|{\sigma^{-1}\left(s,y_{s}\right)}-{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left\|{\sigma^{-1}\left(s,y_{s}\right)}+{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left|v\right|^{2}}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}\left\|{\sigma^{-1}\left(s,y_{s}\right)}\right\|_{2}\left\|{\sigma\left(s,\varphi_{s}\right)-\sigma\left(s,y_{s}\right)}\right\|_{2}\left\|{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}$
$\qquad\times\left\|{\sigma^{-1}\left(s,y_{s}\right)}+{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left|v\right|^{2}\,{\rm d}s.$
Since both $\sigma^{-1}$ and $v$ are bounded, and $\sigma$ is Lipschitz continuous, while $\varphi_{t}-X(0)\in\mathbb{H}^{1}$, we conclude that there exists a constant $C>0$ such that
$B_{3}\leq C\left\|\sigma(s,\varphi_{s})-\sigma(s,y_{s})\right\|_{2}\leq\epsilon CL\left\|W^{\sigma}\right\|.$
The above inequality implies that
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{3}\right\}\big{|}\|W^{\sigma}\|\leq 1\right)=1$ (35)
for all $c\in\mathbb{R}$.
For the fourth term $B_{4}$, let $v:=J\nabla H(y_{s})-\dot{\varphi}_{s}$, $u:=J\nabla H(\varphi_{s})-\dot{\varphi}_{s}$, then.
$B_{4}$ $=\frac{1}{2}\int_{0}^{1}{\left|\sigma^{-1}(s,\varphi_{s})u\right|^{2}-\left|\sigma^{-1}(s,\varphi_{s})v\right|^{2}}\,{\rm d}s$
$=\frac{1}{2}\int_{0}^{1}{\left\langle\sigma^{-1}(s,\varphi_{s})^{2}\left(u-v\right),\left(u+v\right)\right\rangle}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left\|\sigma^{-1}(s,\varphi_{s})^{2}\right\|_{2}\left|\left(u-v\right)\right|\left|\left(u+v\right)\right|}\,{\rm d}s.$
Since both $\sigma^{-1}$, $v$, and $u$ are bounded, and $J\nabla H(y_{s})$ is Lipschitz continuous, and noting that $\varphi_{t}-X(0)\in\mathbb{H}^{1}$, we can conclude that there exists a constant $C>0$ such that
$B_{4}\leq C\left|u-v\right|\leq C\left|J\nabla H(\varphi_{s})-J\nabla H(y_{s})\right|\leq\epsilon CL\left\|W^{\sigma}\right\|.$
The above inequality implies that
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{4}\right\}\big{|}\|W^{\sigma}\|\leq 1\right)=1$ (36)
for all $c\in\mathbb{R}$.
[…]
On one hand, we estimate the upper bound of this probability. Suppose that $F$ is a closed set, and under the guarantee of the continuity of the function $U$, we can envision the existence of a sequence of compact sets $\{K_{n}\}$ satisfying $K_{n}\subset F$ and such that $\inf_{\varphi\in K_{n}}U(\varphi)$ converges to $\inf_{\varphi\in F}U(\varphi)$.
For each compact set $K_{n}$ and any $\epsilon>0$, there always exist a finite number of continuous curves $\{\varphi_{j}\}_{j=1}^{N}\subset K_{n}$ and their corresponding neighborhoods $K(\varphi_{j},\epsilon)$ such that
$K_{n}\subset\bigcup_{j=1}^{N}K(\varphi_{j},\epsilon).$
Applying the union bound principle to this finite covering, we obtain
$\mathbb{P}((X(t)\in K_{n})\leq\sum_{j=1}^{N}\mathbb{P}((X(t)\in K(\varphi_{j},\epsilon)).$
Taking the logarithm of the above inequality and multiplying by $\epsilon^{2}$, we get
$\epsilon^{2}\ln\mathbb{P}((X(t)\in K_{n})$ $\leq\epsilon^{2}\ln\left(N\cdot\max_{1\leq j\leq N}\mathbb{P}((X(t)\in K(\varphi_{j},\epsilon))\right)$
$\leq\epsilon^{2}\ln N+\max_{1\leq j\leq N}\epsilon^{2}\ln\mathbb{P}((X(t)\in K(\varphi_{j},\epsilon)).$
Taking the upper limit $\limsup\limits_{\epsilon\to 0}$ of the above equation, we obtain
$\limsup_{\epsilon\to 0}\epsilon^{2}\ln\mathbb{P}((X(t)\in K_{n})\leq\max_{1\leq j\leq N}\left(\limsup_{\epsilon\to 0}\epsilon^{2}\ln\mathbb{P}((X(t)\in K(\varphi_{j},\epsilon))\right).$
Since $\epsilon^{2}\ln N$ tends to 0, the dominant term is determined by the neighborhood with the maximum probability. Combining the above proofs, we have
$\quad\limsup_{\epsilon\to 0}\epsilon^{2}\ln\mathbb{P}((X(t)\in K_{n})$
$\leq\lim_{\epsilon\to 0}\max_{1\leq j\leq N}\left(-\frac{1}{2}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s+\mathbb{E}\left(\sum_{i=1}^{6}B_{i}\mid\left\|W^{\sigma}\right\|\leq 1\right)+\epsilon^{2}\ln\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1)\right)$
$\leq\max_{1\leq j\leq N}\left(-\frac{1}{2}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s\right)$
$\leq-\min_{1\leq j\leq N}\frac{1}{2}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s.$
Since $\{\varphi_{j}\}\subset K_{n}$, we get
$\min_{1\leq j\leq N}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s\geq\inf_{\varphi\in K_{n}}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s.$
Therefore
$\limsup_{\epsilon\to 0}\epsilon^{2}\ln\mathbb{P}((X(t)\in K_{n})\leq-\inf_{\varphi\in K_{n}}\frac{1}{2}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s.$
For any $\delta>0$, there exists a compact set $K_{n}\subset F$ such that
$\inf_{\varphi\in F}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s+\delta\geq\inf_{\varphi\in K_{n}}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s\geq\inf_{\varphi\in F}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s-\delta.$
Since $\mathbb{P}((X(t)\in F)\leq\mathbb{P}((X(t)\in K_{n})+\mathbb{P}((X(t)\in F\setminus K_{n})$, and the probability of the part outside the compact set decays exponentially to 0 as $n$ increases, the main contribution comes from $K_{n}$. Therefore, as $n$ tends to infinity and $\epsilon$ tends to 0, we obtain
$\limsup_{\epsilon\to 0}\epsilon^{2}\ln\mathbb{P}((X(t)\in F)\leq-\inf_{\varphi\in F}\frac{1}{2}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s.$ (37)
On the other hand, we estimate the lower bound of this probability. Let $G$ be an arbitrary open set in $\mathbb{R}^{n}$. For any $\varphi\in G$, it holds that
$\lim_{{\epsilon\to 0}}\mathbb{P}((X(t)\in G)\geq\lim_{{\epsilon\to 0}}\mathbb{P}(X^{\epsilon}\in K(\varphi,\epsilon)).$
Since $\varphi\in G$ is arbitrarily chosen, we have
$\lim_{{\epsilon\to 0}}\mathbb{P}((X(t)\in G)\geq\lim_{\epsilon\to 0}\sup_{\varphi\in G}\mathbb{P}(X^{\epsilon}\in K(\varphi,\epsilon)).$
That is,
$\liminf_{{\epsilon\to 0}}\epsilon^{2}\ln\mathbb{P}(X(t)\in G)\geq-\inf_{\varphi\in G}\frac{1}{2}\int_{0}^{1}U(\varphi)\,{\rm d}s.$ (38)
By combining the upper bound inequality (37) and the lower bound inequality (38), we can derive the rate function for the stochastic Hamiltonian system
$S(\varphi)=\frac{1}{2}\left(\int_{0}^{1}\left\|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right\|^{2}\,{\rm d}s\right).$
6 Preservation of Invariant Tori in Nearly Integrable Stochastic Hamiltonian Systems
In this section, we investigate the Onsager-Machlup functional, most probable paths, and large deviation principles in nearly integrable Hamiltonian systems with stochastic perturbations. By integrating these findings with KAM theory, we further examine the persistence of invariant tori in almost integrable Hamiltonian systems under small stochastic perturbations, particularly in the most probable sense. To apply KAM theory, in classical mechanics, converting generalized coordinates and momenta into action-angle variables greatly simplifies the analysis of integrable systems.
Firstly, we introduce some notations that will be used throughout this section:
  • The $d$-dimensional torus is denoted by $\mathbb{T}^{d}:=\mathbb{R}^{d}/2\pi\mathbb{Z}^{d}$.
  • For $\alpha>0$ and $\tau\geq d-1\geq 1$, the set of $(\alpha,\tau)$-Diophantine numbers in $\mathbb{R}^{d}$ is defined as
    $\Delta_{\alpha}^{\tau}:=\left\{\omega\in\mathbb{R}^{d}:|\omega\cdot k|\geq\frac{\alpha}{|k|_{1}^{\tau}},\forall 0\neq k\in\mathbb{Z}^{d}\right\}.$
  • The Lebesgue (outer) measure on $\mathbb{R}^{d}$ is denoted by measmeas.
  • For $l\in\mathbb{R}$, the integer part is denoted by $[l]$ and the fractional part by $\{l\}$.
  • For $l>0$ and an open subset $A$ of $\mathbb{R}^{d}$ or $\mathbb{R}^{d}\times\mathbb{T}^{d}$, the set $C^{l}(A)$ consists of continuously differentiable functions $f$ on $A$ up to the order $[l]$ such that $f^{[l]}$ is Hölder-continuous with exponent $\{l\}$ and with finite $C^{l}$-norm defined by
    $\|f\|_{C^{l}(A)}$ $=\max\left\{\|f\|_{C^{[l]}(A)},\|f^{[l]}\|_{C^{\{l\}}(A)}\right\},$
    $\|f\|_{C^{[l]}(A)}$ $=\max_{\begin{subarray}{c}k\in\mathbb{N}^{d}\\ 0\leq|k|_{1}\leq[l]\end{subarray}}\sup_{A}|\partial_{I}^{k}f|,$
    $\|f^{[l]}\|_{C^{\{l\}}(A)}$ $=\max_{\begin{subarray}{c}k\in\mathbb{N}^{d}\\ |k|_{1}=[l]\end{subarray}}\sup_{\begin{subarray}{c}I_{1},I_{2}\in A\\ 0<|I_{1}-I_{2}|<1\end{subarray}}\frac{|\partial_{I}^{k}f(I_{1})-\partial_{I}^{k}f(I_{2})|}{|I_{1}-I_{2}|^{\{l\}}}.$
    When $A=\mathbb{R}^{d}$ or $A=\mathbb{R}^{d}\times\mathbb{T}^{d}$, we simplify this to $\|f\|_{C^{l}}$.
  • For $l>0$ and any subset $A$ of $\mathbb{R}^{d}$, $C_{W}^{l}(A)$ denotes the set of functions of class $C^{l}$ on $A$ in the sense of Whitney.
  • For $r,s>0$, $y_{0}\in\mathbb{C}^{d}$, and $\varnothing\neq\mathscr{D}\subseteq\mathbb{C}^{d}$, we define
    $\begin{split}\mathbb{T}_{s}^{d}&:=\left\{x\in\mathbb{C}^{d}:|\text{Im }x|<s\right\}/2\pi\mathbb{Z}^{d},\\ B_{r}(y_{0})&:=\left\{y\in\mathbb{R}^{d}:|y-y_{0}|<r\right\},\quad(y_{0}\in\mathbb{R}^{d}),\\ D_{r}(y_{0})&:=\left\{y\in\mathbb{C}^{d}:|y-y_{0}|<r\right\},\\ D_{r,s}(y_{0})&:=D_{r}(y_{0})\times\mathbb{T}_{s}^{d},\\ D_{r,s}(\mathscr{D})&:=\bigcup_{y_{0}\in\mathscr{D}}D_{r,s}(y_{0}).\end{split}$
  • The unit $(d\times d)$ matrix is denoted by $\mathbb{I}_{d}:=\text{diag}(1)$, and the standard symplectic matrix is given by
    $\mathbb{J}:=\begin{pmatrix}0&-\mathbb{I}_{d}\\ \mathbb{I}_{d}&0\end{pmatrix}$.
  • For $\mathscr{D}\subseteq\mathbb{C}^{d}$, $A_{r,s}(\mathscr{D})$ denotes the Banach space of real-analytic functions with bounded holomorphic extensions to $D_{r,s}(\mathscr{D})$, with norm
    $\|\cdot\|_{r,s,\mathscr{D}}:=\sup_{D_{r,s}(\mathscr{D})}|\cdot|.$
  • The canonical symplectic form on $\mathbb{C}^{d}\times\mathbb{C}^{d}$ is given by
    $\omega:=d\theta\land dI=d\theta_{1}\land dI_{1}+\cdots+d\theta_{d}\land dI_{d},$
    and $\phi_{H}^{t}$ denotes the associated Hamiltonian flow governed by the Hamiltonian $H(I,\theta)$, $I,\theta\in\mathbb{C}^{d}$.
  • The projections on the first and last $d$-components are denoted by $\pi_{1}:\mathbb{C}^{d}\times\mathbb{C}^{d}\ni(I,\theta)\longmapsto I$ and $\pi_{2}:\mathbb{C}^{d}\times\mathbb{C}^{d}\ni(I,\theta)\longmapsto\theta$, respectively.
  • For a linear operator $\mathcal{L}$ from the normed space $(V_{1},\|\cdot\|_{1})$ into the normed space $(V_{2},\|\cdot\|_{2})$, its ”operator norm” is given by
    $\|\mathcal{L}\|:=\sup_{x\in V_{1}\setminus\{0\}}\frac{\|\mathcal{L}x\|_{2}}{\|x\|_{1}},$
    so that $\|\mathcal{L}x\|_{2}\leq\|\mathcal{L}\|\|x\|_{1}$ for any $x\in V_{1}$.
  • For $\omega\in\mathbb{R}^{d}$ and a $C^{1}$ function $f$, the directional derivative of $f$ with respect to $\omega$ is given by
    $D_{\omega}f:=\omega\cdot f_{I}=\sum_{j=1}^{d}\omega_{j}f_{I_{j}}.$
  • If $f$ is a smooth or analytic function on $\mathbb{T}^{d}$, its Fourier expansion is given by
    $f=\sum_{k\in\mathbb{Z}^{d}}f_{k}e^{ik\cdot\theta},\quad f_{k}:=\frac{1}{(2\pi)^{d}}\int_{\mathbb{T}^{d}}f(\theta)e^{-ik\cdot\theta}d\theta,$
    where $e:=\exp(1)$ denotes the Neper number and $i$ the imaginary unit. We also set
    $\langle f\rangle:=f_{0}=\frac{1}{(2\pi)^{d}}\int_{\mathbb{T}^{d}}f(\theta)d\theta.$
Without loss of generality, the Hamiltonian equations for an integrable system are expressed as
$\begin{cases}d\theta(t)=\frac{\partial H}{\partial I}(I(t))\,dt=\omega(I(t))\,dt,\\ dI(t)=0,\end{cases}$ (39)
where $I$ are action variables, $\theta$ are angle variables, and $\omega(I)$ are frequencies associated with the action variables. For an integrable system, the action variables $I$ remain constant over time, while the angle variables $\theta$ evolve linearly with time.
When introducing a small deterministic perturbation $\epsilon_{1}P(I,\theta)$, the almost integrable Hamiltonian becomes
$H_{\epsilon}(I,\theta)=H(I)+\epsilon_{1}P(I,\theta),$
and the corresponding Hamiltonian equations in action-angle variables take the following form:
$\mathrm{d}X(t)=J\nabla H_{\epsilon}(X(t))\,\mathrm{d}t.$ (40)
In this scenario, the action variables $I$ are no longer constant and undergo slight changes due to the perturbation, while the evolution of the angle variables $\theta$ is correspondingly modified.
When a small stochastic perturbation $\epsilon_{2}\sigma(t)\,dW(t)$ is further introduced, the corresponding stochastic Hamiltonian system in action-angle variables is described by the following stochastic differential equations:
$\mathrm{d}X(t)=J\nabla H_{\epsilon}(X(t))\,\mathrm{d}t+\epsilon_{2}\sigma(t,X(t))\,\mathrm{d}W(t),$ (41)
where $\epsilon_{2}$ represents the strength of the stochastic perturbation, $\sigma(t,X(t))$ denotes the diffusion coefficient matrix, and $W(t)$ is a $2n$-dimensional Wiener process with independent components.
We then investigate the invariant tori associated with the Hamiltonian $H_{\epsilon}(I,\theta)$ corresponding to Equation (40). The following assumptions are imposed:
  1. (H1)
    Let $l>2\nu:=2(\tau+1)>2d\geq 4$, and let $\mathscr{D}\subset\mathbb{R}^{d}$ be a non - empty, bounded domain.
  2. (H2)
    Consider the Hamiltonian $H_{\epsilon}(I,\theta)$ on the phase space $\mathscr{D}\times\mathbb{T}^{d}$. Here, $H$ and $P$ are given functions in $C^{l}(\mathscr{D}\times\mathbb{T}^{d})$ with finite $l$ - norms $\|H\|_{C^{l}(\mathscr{D})}$ and $\|P\|_{C^{l}(\mathscr{D})}$.
  3. (H3)
    Assume that $H_{I}$ is locally uniformly invertible. This implies that for all $I\in\mathscr{D}$, $\det H_{II}(I)\neq 0$. To simplify notation, define $T(I):=H_{II}(I)^{-1}$, and suppose that $C_{T}:=\|T\|_{C^{0}(\mathscr{D})}<\infty$. Furthermore, set $C_{H}:=\max\left\{1,\|H\|_{C^{l}(\mathscr{D})}\right\}$, and then define $\theta:=C_{T}C_{H}$, which possesses the property that $\theta\geq 1$.
  4. (H4)
    Let $\alpha\in(0,1)$ and set,
    $\quad\alpha_{*}:=\alpha^{\frac{1}{l-2v}},\quad\mathscr{D}^{\prime}:=\{I\in\mathscr{D}:B_{\alpha_{*}}(I)\subseteq\mathscr{D}\},\quad\mathscr{D}_{\alpha}:=\{I\in\mathscr{D}^{\prime}:H_{I}(I)\in\Delta^{\tau}_{\alpha}\}.$
  5. (H5)
    Finally, for some suitable constant $C_{1}=C_{1}(d,l)>1$, set
    $\begin{cases}\sigma:=\left(\frac{\varepsilon^{3/2}}{\theta^{2l/\nu}\alpha\sqrt{C_{H}}}\right)^{1/(l+\nu)},\\ \rho:=\frac{2C_{1}C_{H}\varepsilon}{\alpha^{2}\sigma^{2\nu}},\\ \beta_{0}:=\min\left\{\frac{l}{2\nu}-1+\frac{1}{\nu},2\right\}.\end{cases}$
Under the aforementioned notation and assumptions (C1), (C2), and (H1)-(H5), the following stochastic version of the KAM theorem holds.
Theorem 6.1.
[…]
$OM(\varphi,\dot{\varphi})$ $=\int_{0}^{1}{\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H_{\epsilon}(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s$
$\quad-\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H_{\epsilon}(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s$
Furthermore, by applying the variational principle to minimize the Onsager-Machlup functional, the most probable transition path can be obtained, corresponding to the solution of the nearly integrable Hamiltonian system
$\dot{\varphi}(t)=J\nabla H_{\epsilon}(\varphi_{t})-\frac{1}{2}\epsilon_{2}^{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}.$
[…]
$\epsilon_{2}^{2}\ln\mathbb{P}(X_{\epsilon}(t)\in A)\approx-\inf_{\varphi\in A}S(\varphi),$
where $A\in\mathbb{R}^{d}$ denotes an arbitrary measurable set and the rate function $S(\varphi)$ is given by
$S(\varphi)=\frac{1}{2}\left(\int_{0}^{1}\left\|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H_{\epsilon}(\varphi_{s})-\dot{\varphi_{s}}\right)\right\|^{2}\,{\rm d}s\right).$
Proof.
[…]
Next, we will outline the framework for proving the existence of invariant tori and the measure of Cantor sets, with detailed proofs referred to in [26]. This commences by extending $H$ and $P$ to the entirety of the phase space $\mathbb{R}^{d}\times\mathbb{T}^{d}$. To accomplish this extension, we introduce a cut-off function $\chi\in C(\mathbb{C}^{d})\cap C^{\infty}(\mathbb{R}^{d})$ that fulfills the conditions $0\leq\chi\leq 1$, with its support confined within $D_{\alpha_{*}}(\mathscr{D}^{\prime})$ and $\chi$ being identically equal to 1 on $D_{\alpha_{*}/2}(\mathscr{D}^{\prime})$. Additionally, for any multi-index $k\in\mathbb{N}^{d}$ with $|k|_{1}\leq l$, there exists a constant $C_{0}=C_{0}(d,l)>0$ such that
$\|\partial_{y}^{k}\chi\|_{\mathbb{R}^{d}}\leq C_{0}\alpha_{*}^{-|k|_{1}}.$
Utilizing Faà Di Bruno’s Formula [7], we construct $\widehat{H}\in C^{l}(\mathbb{R}^{d})$ such that $\|T\|_{\mathscr{D}}\|\widehat{H}-H\|_{C^{l}(\mathscr{D})}\leq C_{1}^{-1}\alpha_{*}^{l}/4$. Consequently, $\widehat{H}_{II}=H_{II}(\mathbb{I}_{d}+T(\widehat{H}_{II}-H_{II}))$ is invertible on $\mathscr{D}$ with $\|(\widehat{H}_{II})^{-1}\|_{\mathscr{D}}\leq 2\|T\|_{\mathscr{D}}$.
[…]
$\|\tilde{H}\|_{C^{l}}\leq\|H\|_{C^{l}}+C_{1}\alpha_{*}^{-l}\|\widehat{H}-H\|_{C^{l}}<2\|H\|_{C^{l}},$
and
$\|(\widehat{H}_{II})^{-1}\partial_{y}^{2}(\chi\cdot(H-\widehat{H}))\|_{\mathscr{D}}\leq\|(\widehat{H}_{II})^{-1}\|_{\mathscr{D}}\cdot C_{1}\alpha_{*}^{-l}\|\widehat{H}-H\|_{C^{l}}\leq 1/2.$
Thus, $\tilde{H}_{II}$ is invertible with $\|(\tilde{H}_{II})^{-1}\|_{\mathscr{D}}\leq 2\|(\widehat{H}_{II})^{-1}\|_{\mathscr{D}}\leq 4\|T\|_{\mathscr{D}}$.
Similarly, $P$ is extended to a function $\tilde{P}\in C^{l}(\mathbb{R}^{d}\times\mathbb{T}^{d})$ such that $\tilde{P}\equiv P$ on $D_{\alpha_{*}/2}(\mathscr{D}^{\prime})$ and $\|P\|_{C^{l}}\leq 2\|P\|_{C^{l}}$. Setting $\tilde{H}_{\epsilon}:=\tilde{H}+\tilde{P}$, we observe $\tilde{H}_{\epsilon}|_{D_{\alpha_{*}/2}(\mathscr{D}^{\prime})}={H}_{\epsilon}$. Notably, replacing ${H}_{\epsilon}$ with $\tilde{H}_{\epsilon}$ makes no difference since the invariant tori of $\tilde{H}_{\epsilon}$ we aim to construct reside within $D_{\alpha_{*}/2}(\mathscr{D}^{\prime})$, given $r_{0}<\alpha_{*}/2$.
Let $\mathcal{D}\subset\mathbb{R}^{d}$ be a domain with a smooth boundary $\partial\mathcal{D}$. Suppose there exists a positive constant $c=c(d,\tau,l)<1$ such that the parameters $\alpha$ and $\varepsilon$ satisfy the following conditions:
$0<\alpha\leq\min\left\{cC_{H},\frac{R(\mathcal{D})}{6},\frac{1}{2}\,\mathrm{minfoc}(\partial\mathcal{D})\right\},\quad\varepsilon\leq cC_{H}^{-\frac{l+2v}{l-2v}}\theta^{-a}\alpha^{\frac{2l}{l-2v}},$
[…]
Let $\mathcal{H}_{j}$ (resp. $\mathcal{P}_{j}$) denote the real-analytic approximation $H_{\xi_{j}}$ (resp. $P_{\xi_{j}}$) of $\tilde{H}$ (resp. $\tilde{P}$) defined on $\mathcal{O}_{j}$, as given by Lemma 2.17. Define the initial set $\mathscr{D}_{0}$ as
$\mathscr{D}_{0}:=\{I\in\mathbb{R}^{d}:\partial_{I}\mathcal{H}_{0}(I)\in\partial_{I}H(\mathscr{D}_{\alpha})\}.$
For $j\geq 1$, we define the proposition $(\mathscr{P}_{j})$ as follows: There exist:
  1. 1.
    A sequence of sets $\mathscr{D}_{j}$,
  2. 2.
    A sequence of diffeomorphisms $G_{j}:D_{\tilde{r}_{j}}(\mathscr{D}_{j-1})\to G_{j}(D_{\tilde{r}_{j}}(\mathscr{D}_{j-1}))$,
  3. 3.
    A sequence of real-analytic symplectic transformations
    $\Phi_{j}=(v_{j},u_{j}):D_{r_{j},s_{j}}(\mathscr{D}_{j})\to D_{\sigma_{j-1},\sigma_{j-1}}(\mathscr{D}_{j-1}),$
such that, setting $\mathcal{H}^{\epsilon}_{j-1}:=\mathcal{H}_{j-1}+\mathcal{P}_{j-1}$, the following properties hold:
$G_{j}(\mathscr{D}_{j-1})=\mathscr{D}_{j}\subset\mathscr{D}_{r_{j}},\quad G_{j}=(\partial_{I}H_{j})^{-1}\circ\partial_{I}H_{j-1},$
$\det\partial_{I}^{2}H_{j}(y)\neq 0,\quad T_{j}(I):=\partial_{I}^{2}H_{j}(I)^{-1},\quad\forall I\in\mathscr{D}_{j},$
$H^{\epsilon}_{j}:=\mathcal{H}^{\epsilon}_{j-1}\circ\phi^{j}:=H_{j}+P_{j}\quad\text{on }D_{r_{j},s_{j}}(\mathscr{D}_{j}),$
where $\phi^{j}:=\phi_{1}\circ\phi_{2}\circ\cdots\circ\phi_{j}$ and $K_{0}:=\mathcal{K}_{0}$.
Moreover,
$\left\lVert G_{j}-\text{id}\right\rVert_{\tilde{r}_{j},\mathscr{D}_{j-1}}\leq\tilde{r}_{j}\xi^{2\nu}\xi^{m(j-1)},\quad\left\lVert\partial_{y}G_{j}-\mathbb{I}_{d}\right\rVert_{\tilde{r}_{j},\mathscr{D}_{j-1}}\leq\xi^{2\nu}\xi^{2m(j-1)},$
$\left\lVert\partial_{I}^{2}H_{j}\right\rVert_{r_{j},\mathscr{D}_{j}}<2C_{H},\quad\left\lVert T_{j}\right\rVert_{\mathscr{D}_{j}}<2C_{T},$
$T_{j}:=(\partial_{I}^{2}H_{j})^{-1},\quad\left\lVert P_{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}\leq C_{1}C_{H}\xi_{j}^{l-1},$
$\max\left\{\left\lVert M_{j}(\phi_{j}-\text{id})\right\rVert_{2r_{j},s_{j},\mathscr{D}_{j}},\left\lVert\pi_{2}\partial_{x}(\phi_{j}-\text{id})\right\rVert_{2r_{j},s_{j},\mathscr{D}_{j}}\right\}\leq\xi^{2\nu}\xi^{m(j-1)},$
where $M_{j}:=\text{diag}(r_{j}^{-1}\mathbb{I}_{d},\sigma_{j}^{-1}\mathbb{I}_{d})$.
We will use mathematical induction to prove that the proposition $(\mathscr{P}_{j})$ holds for all $j\geq 1$. The proof of this part primarily relies on the application of Lemma 2.16. It is straightforward to verify that Lemma 2.16 can be applied to $\mathcal{H}^{\epsilon}_{0}$, which implies that the statement $(\mathscr{P}_{1})$ holds.
Next, we assume that $(\mathscr{P}_{j})$ holds for some $j\geq 1$ and proceed to prove $(\mathscr{P}_{j+1})$. First, observe the following estimates:
$s_{j}+\frac{\sigma_{j-1}}{3}=\frac{(12\xi+1)\sigma_{j-1}}{3}<\frac{2\sigma_{j-1}}{3}<\frac{s_{j-1}}{2},$
$2r_{j}+\frac{r_{j-1}\sigma_{j-1}}{3}<\frac{r_{j-1}}{4}+\frac{r_{j-1}}{6}<\frac{r_{j-1}}{2},$
$2r_{j}+\frac{r_{j-1}\sigma_{j-1}}{3}<\sigma_{0}^{\nu}\xi^{j}+\frac{\sigma_{j-1}}{3}=\sigma_{0}^{\tau}\sigma_{j}+\frac{\sigma_{j-1}}{3}<\sigma_{j-1}.$
These inequalities, combined with a symplectic change of coordinates $\phi^{\prime}=\text{id}+\tilde{\phi}:D_{\overline{r}/2,s^{\prime}}(\mathscr{D}^{\prime})\to D_{\overline{r}+r\sigma/3,\overline{s}}(\mathscr{D})$ in Lemma 2.16, imply that
$\phi_{j}(D_{r_{j},s_{j}}(\mathscr{D}_{j}))\subset D_{\sigma_{j-1},\sigma_{j-1}}(\mathscr{D}_{j-1})\bigcap D_{r_{j-1}/2,s_{j-1}/2}(\mathscr{D}_{j-1}).$ (42)
In particular, the real-analytic symplectic transformation
$\phi_{j}=(v_{j},u_{j}):D_{r_{j},s_{j}}(\mathscr{D}_{j})\to D_{\sigma_{j-1},\sigma_{j-1}}(\mathscr{D}_{j-1})$
exists. Furthermore, the inequality
$2r_{j+1}<\frac{1}{4}\min\left\{\frac{\alpha}{2d(2C_{H})\kappa_{j}^{\nu}},\check{r}_{j+1}\right\}$
holds. Together with the definitions of the sequences of the various parameters, this ensures that condition (4) in Lemma 2.16 is satisfied for all $j\geq 1$.
Write
$\mathcal{H}^{\epsilon}_{j}:=\mathcal{H}_{j}+\mathcal{P}_{j}=\mathcal{H}^{\epsilon}_{j-1}+(\mathcal{H}_{j}-\mathcal{H}_{j-1})+(\mathcal{P}_{j}-\mathcal{P}_{j-1}).$
By the inductive assumption and (42), we have
$\mathcal{H}^{\epsilon}_{j}\circ\phi^{j}$ $=\mathcal{H}^{\epsilon}_{j-1}\circ\phi^{j}+(\mathcal{H}_{j}-\mathcal{H}_{j-1})\circ v^{j}+(\mathcal{P}_{j}-\mathcal{P}_{j-1})\circ\phi^{j}$
$=H_{j}+P_{j}+(\mathcal{H}_{j}-\mathcal{H}_{j-1})\circ v_{j}+(\mathcal{P}_{j}-\mathcal{P}_{j-1})\circ\phi^{j}$
$=\mathcal{H}^{j}+\mathcal{P}^{j},\quad\text{on }D_{r_{j},s_{j}}(\mathscr{D}_{j}),$
where $\mathcal{H}^{j}:=H_{j}$ and $\mathcal{P}^{j}:=P_{j}+(\mathcal{H}_{j}-\mathcal{H}_{j-1})\circ\phi^{j}+(\mathcal{P}_{j}-\mathcal{P}_{j-1})\circ\phi^{j}$, with
$\left\lVert\partial_{y}^{2}\mathcal{H}^{j}\right\rVert_{r_{j},\mathscr{D}_{j}}<2C_{K},\qquad\left\lVert(\partial_{y}^{2}\mathcal{H}^{j})^{-1}\right\rVert_{\mathscr{D}_{j}}<2C_{T},$ (43)
by the inductive assumption, provided $\phi^{j}$ maps $D_{r_{j},s_{j}}(\mathscr{D}_{j})$ into $\mathcal{O}_{j}=\{(I,\theta)\in\mathbb{C}^{d}\times\mathbb{C}^{d}:|\text{Im}(I,\theta)|<\xi_{j}\}$ i.e.
$\sup_{D_{r_{j},s_{j}}(\mathscr{D}_{j})}|\text{Im}\phi^{j}|\leq\frac{\xi_{j}}{2}.$ (44)
Hence
$\left\lVert\mathcal{P}^{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}$ $\leq\left\lVert P_{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}+\left\lVert(\mathcal{H}_{j}-\mathcal{H}_{j-1})\circ\phi^{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}+\left\lVert(\mathcal{P}_{j}-\mathcal{P}_{j-1})\circ\phi^{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}$
$\leq C_{1}C_{H}\xi_{j-1}^{l}+\left\lVert\mathcal{H}_{j}-\mathcal{H}_{j-1}\right\rVert_{\xi_{j}}+\left\lVert\mathcal{P}_{j}-\mathcal{P}_{j-1}\right\rVert_{\xi_{j}}$
$\leq C_{1}C_{H}\xi_{j-1}^{l}+C_{1}C_{H}\xi_{j-1}^{l}+C_{1}\varepsilon\xi_{j-1}^{l}$
$\leq 3C_{1}C_{H}\xi_{j-1}^{l}.$
Thus, thanks to (43), $\mathcal{H}^{\epsilon}_{j}\circ\phi^{j}=\mathcal{H}^{j}+\mathcal{P}^{j}$ satisfies the assumptions in (3) with $\varepsilon\sim\left\lVert\mathcal{P}^{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}$, $r\sim r_{j}$, $s\sim s_{j}$, $\sigma\sim\sigma_{j}$, $C_{H}\sim 2C_{H}$ as
$\partial_{I}^{2}\mathcal{H}^{j}(\mathscr{D}_{j})\stackrel{{\scriptstyle\text{def}}}{{=}}\partial_{I}^{2}H_{j}(G_{j}(\mathscr{D}_{j-1}))=\partial_{I}^{2}H_{j-1}(\mathscr{D}_{j-1})=\cdots=\partial_{I}^{2}H_{0}(\mathscr{D}_{0})\subset\Delta_{\tau}^{\alpha}.$
Hence, in order to apply Lemma 6 to $\mathcal{H}^{\epsilon}_{j}\circ\phi^{j}=\mathcal{H}^{j}+\mathcal{P}^{j}$, we only need to check (5). Upon observation, we have
$r_{j}=r_{0}\xi^{\nu j}\leq\frac{\alpha}{2C_{H}}\sigma_{j}^{\nu}\leq\alpha\sigma_{j}^{\nu}/\left\lVert\partial_{I}^{2}\mathcal{H}^{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}},$
$\sigma_{1}^{-\nu}\frac{\left\lVert\mathcal{P}^{1}\right\rVert_{r_{1},s_{1},\mathscr{D}_{1}}}{\alpha r_{1}}\rho^{-1}\leq C_{2}\sigma_{0}^{l}\frac{C_{H}}{\varepsilon\xi^{2\nu}}\leq 1,$
$3C_{0}\frac{\theta C_{T}\left\lVert\mathcal{P}^{1}\right\rVert_{r_{1},s_{1},\mathscr{D}_{1}}}{r_{1}\check{r}_{2}\bar{\sigma}_{1}}\leq C_{2}\sigma_{0}^{l-2\nu}\frac{\theta^{6+m}(C_{H})^{2}}{\alpha^{2}}\lambda^{2(\nu+m)}\leq\xi^{2\nu},$
and, for $j\geq 2$,
$\sigma_{j}^{-\nu}\frac{\left\lVert\mathcal{P}^{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}}{\alpha r_{j}}\rho^{-1}\leq C_{2}\sigma_{0}^{l}\frac{C_{H}}{\varepsilon}\xi^{(l-2\nu)j-2l}\leq C_{2}\sigma_{0}^{l}\frac{C_{H}}{\varepsilon}\xi^{-2l}\leq 1,$
$3C_{0}\frac{\theta C_{T}\left\lVert\mathcal{P}^{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}}{r_{j}\check{r}_{j+1}\bar{\sigma}_{j}}\leq C_{2}\sigma_{0}^{l-2\nu}\frac{\theta^{4+2l/\nu}(C_{H})^{2}}{\alpha^{2}}\lambda^{2l}\leq\xi^{2\nu}.$
Therefore, Lemma 6 applies to $\mathcal{H}^{\epsilon}_{j}$ and yields the desired symplectic change of coordinates $\phi_{j+1}$.
Furthermore, based on Lemmas 2.17 and 2.18, we can obtain the convergence results for $G^{j}$, $P_{j}$, $\phi^{j}$, and $H_{j}$ as follows:
  • The sequence $G^{j}:=G_{j}\circ G_{j-1}\circ\cdots\circ G_{2}\circ G_{1}$ converges uniformly on $\mathscr{D}_{0}$ to a diffeomorphism $G_{*}:\mathscr{D}_{0}\to\mathscr{D}_{*}:=G_{*}(\mathscr{D}_{0})\subset\mathscr{D}$, and $G_{*}\in C_{W}^{1}(\mathscr{D}_{0})$;
  • $P_{j}$ converges uniformly to 0 on $\mathscr{D}_{*}\times\mathbb{T}_{s_{*}}^{d}$ in the $C_{W}^{2}$ topology;
  • $\phi^{j}$ converges uniformly on $\mathscr{D}_{*}\times\mathbb{T}^{d}$ to a symplectic transformation
    $\phi_{*}:\mathscr{D}_{*}\times\mathbb{T}^{d}\stackrel{{\scriptstyle\text{into}}}{{\longrightarrow}}\mathscr{D}\times\mathbb{T}^{d},$
    with $\phi_{*}\in C_{W}^{\tilde{m}}(\mathscr{D}_{*}\times\mathbb{T}^{d})$ and $\phi_{*}(\cdot,\cdot)\in C^{\tilde{m}v}(\mathbb{T}^{d})$, for any given $y_{*}\in\mathscr{D}_{*}$;
  • $H_{j}$ converges uniformly on $\mathscr{D}_{*}$ to a function $H_{*}\in C_{W}^{2+\tilde{m}}(\mathscr{D}_{*})$, with
    $\partial_{I_{*}}H_{*}\circ G_{*}=\partial_{I}\mathcal{H}_{0},\qquad\qquad\qquad\text{on }\mathscr{D}_{0},$
    $H\circ\phi_{*}(I_{*},x)=H_{*}(I_{*}),\qquad\forall(I_{*},x)\in\mathscr{D}_{*}\times\mathbb{T}^{d}.$
Based on the preceding proof, we can demonstrate that there exist a Cantor-like set $\mathcal{D}_{*}\subset\mathcal{D}$, an embedding $\phi_{*}=(v_{*},u_{*}):\mathcal{D}_{*}\times\mathbb{T}^{d}\rightarrow\mathcal{H}:=\phi_{*}(\mathcal{D}_{*}\times\mathbb{T}^{d})\subset\mathcal{D}\times\mathbb{T}^{d}$ of class $C_{W}^{\beta}(\mathcal{D}_{*}\times\mathbb{T}^{d})$, and a function $H_{*}\in C_{W}^{2}(\mathcal{D}_{*},\mathbb{R})$, such that $H^{\epsilon}\circ\phi_{*}(I_{*},\theta)=K_{*}(I_{*},\theta)$ for all $(I_{*},\theta)\in\mathcal{D}_{*}\times\mathbb{T}^{d}$. The map $\xi\mapsto\phi_{*}(I_{*},\theta)$ is of class $C_{\beta}^{v}(\mathbb{T}^{d})$ for any $I_{*}\in\mathcal{D}_{*}$ (with $v^{-1}<\beta<\beta_{0}$), and the map $G^{*}:=(\partial_{I_{*}}H_{*})^{-1}\circ\partial_{I}H:\mathcal{D}_{*}\rightarrow\mathcal{D}_{*}$ defines a lipeomorphism onto $\mathcal{D}_{*}$, satisfying $B_{\alpha/2}(\mathcal{D}_{*})\subseteq\mathcal{D}$. The set $\mathcal{K}$ is foliated by KAM tori of $H^{\epsilon}$, each being a graph of a $C^{v}(\mathbb{T}^{d})$-map.
Furthermore, the following estimates hold:
$\|G^{*}-\mathrm{id}\|_{\mathcal{D}_{*}}\leq\varepsilon^{\frac{3\tau}{2(l+v)}}\alpha^{\frac{l+1}{l+v}}C_{H}^{\frac{\tau}{2(l+v)}}\theta^{-1}L^{\frac{2l\tau}{v(l+v)}},\quad\|G^{*}-\mathrm{id}\|_{L,\mathcal{D}_{*}}<\frac{1}{2},$
and
$\sup_{\mathcal{D}_{*}\times\mathbb{T}^{d}}\max\left\{|M(\phi_{*}-\mathrm{id})|,\,\|\pi_{2}(\partial_{x}\phi_{*}-\mathbb{I}_{d})\|\right\}\leq 8\theta^{-2}(\log\rho^{-1})^{-2v}<1,$
where $M:=\mathrm{diag}\left(C_{H}(\alpha\sigma^{v})^{-1}\mathbb{I}_{d},\sigma^{-1}\mathbb{I}_{d}\right)$.
Then, the measure of the complement of $\mathcal{K}$ is bounded by
$\mathrm{meas}(\mathcal{D}\times\mathbb{T}^{d}\setminus\mathcal{K})\leq(3\pi)^{d}\left(2\mathcal{H}^{d-1}(\partial\mathcal{D})\tilde{\epsilon}+C\tilde{\epsilon}^{2}+\mathrm{meas}(\mathcal{D}^{\prime}\setminus\mathcal{D}_{\alpha})\right),$
with
$\tilde{\epsilon}:=\max\left\{\epsilon^{\frac{3\tau}{2(l+v)}}\alpha^{\frac{l+1}{l+v}}C_{H}^{\frac{-\tau}{2(l+v)}}\theta^{-1}L^{\frac{2l\tau}{v(l+v)}},\alpha_{*}\right\},C=2\sum_{j=1}^{\left\lfloor\frac{d-1}{2}\right\rfloor}\frac{\tilde{\epsilon}^{2j-1}{k}_{2j}(\mathbb{R}^{\partial\mathcal{D}})}{1\cdot 3\cdots(2j+1)},$
where ${k}_{2j}(\mathbb{R}^{\partial\mathcal{D}})$ denotes the $(2j)$-th integrated mean curvature of $\partial\mathcal{D}$.
[…]
$\epsilon^{2}\ln\mathbb{P}(X(t)\in A)\approx-\inf_{\varphi\in A}S(\varphi),$
where $\varphi-x_{0}\in\mathbb{H}^{1}$, $A\in\mathbb{R}^{d}$ denotes an arbitrary measurable set and the rate function $S(\varphi)$ is given by
$S(\varphi)$ $=\frac{1}{2}\left(\int_{0}^{1}\left\|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H_{\epsilon}(\varphi_{s})-\dot{\varphi_{s}}\right)\right\|^{2}\,{\rm d}s\right).$
This result aligns with the conclusions of the deterministic KAM theory, further demonstrating that most invariant tori can survive under small perturbations. In a stochastic setting, these tori are preserved in a probabilistic sense, providing new theoretical insights into the stability of nearly integrable Hamiltonian systems under stochastic perturbations. Based on Theorem 6.1, we derive an intriguing corollary.
Corollary 6.2.
[…]
$\mathbb{P}(X_{\epsilon}(t)\in\mathcal{K}_{0})\approx\exp\left\{-C\frac{\epsilon_{1}^{2}}{\epsilon_{2}^{2}}\right\}\times\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1).$
Therefore, when the ratio of the strength of the deterministic perturbation $\epsilon_{1}$ to the strength of the stochastic noise term $\epsilon_{2}$, denoted as $\frac{\epsilon_{1}}{\epsilon_{2}}$, tends to 0, the probability that the solution $X_{\epsilon}(t)$ remains on the invariant torus $\mathcal{K}_{0}$ is equal to $\mathbb{P}(\|W^{\sigma}\|\leq 1)$. Conversely, when $\frac{\epsilon_{1}}{\epsilon_{2}}$ tends to infinity, the probability that the solution $X_{\epsilon}(t)$ remains on the invariant torus $\mathcal{K}_{0}$ is equal to 0.
Proof.
Based on the content of Theorem 6.1, we have
$\quad\frac{\mathbb{P}(X_{\epsilon}(t)\in\mathcal{K}_{0})}{\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1}\approx\exp\left\{-\frac{1}{\epsilon_{2}^{2}}\inf_{\varphi\in\mathcal{K}_{0}}S(\varphi)\right\}$
$=\exp\left\{-\frac{1}{2\epsilon_{2}^{2}}\inf_{\varphi\in\mathcal{K}_{0}}\left(\int_{0}^{1}\left\|\sigma^{-1}(t)\left(\epsilon_{1}J\nabla P(\varphi_{s})\right)\right\|^{2}\,{\rm d}s\right)\right\}$
$=\exp\left\{-\frac{\epsilon_{1}^{2}}{2\epsilon_{2}^{2}}\inf_{\varphi\in\mathcal{K}_{0}}\left(\int_{0}^{1}\left\|\sigma^{-1}(t)\left(\epsilon_{1}J\nabla P(\varphi_{s})\right)\right\|^{2}\,{\rm d}s\right)\right\}$
$=\exp\left\{-C\frac{\epsilon_{1}^{2}}{\epsilon_{2}^{2}}\right\},$
where $C$ is a quantity that depends on $\sigma^{-1}$ and the partial derivatives of $P$. ∎
Example 6.3.
To further illustrate the above theory, we introduce a two-dimensional stochastic oscillator equation as a concrete example. This system describes two coupled harmonic oscillators under both deterministic and stochastic perturbations. The stochastic oscillator equations are given as follows:
[…]
In this system, $\epsilon$ represents the perturbation coefficient, and $W_{1}(t),W_{2}(t),W_{3}(t),W_{4}(t)$ are independent Wiener processes that introduce random perturbations into the system. Here, $q_{1}(t)$ and $q_{2}(t)$ are the generalized coordinates, and $p_{1}(t)$ and $p_{2}(t)$ are the corresponding momenta. The presence of the stochastic terms makes the system non-deterministic, subject to random noise.
When the stochastic terms disappear, the system reduces to the following deterministic Hamiltonian system:
[…]
This is a typical nearly integrable Hamiltonian system, where $\epsilon$ represents a small deterministic perturbation. Without random perturbations, the system exhibits classic harmonic oscillatory behavior, with the relationship between the generalized coordinates and momenta governed by the Hamiltonian.
The Hamiltonian of this system can be written as
$H(q_{1},q_{2},p_{1},p_{2},t)=\frac{p_{1}^{2}}{2}+\frac{p_{2}^{2}}{2}+q_{1}^{2}+\frac{q_{2}^{2}}{2}-\epsilon\left(q_{1}\sin(0.6t)+q_{2}\cos(0.6t)+p_{1}p_{2}\right),$
where $q_{1},q_{2}$ are the generalized coordinates and $p_{1},p_{2}$ are their corresponding momenta. The term $\epsilon q_{1}p_{2}\sin(t)$ represents the perturbation, which introduces coupling between the two oscillators and depends on time $t$. This coupling alters the energy distribution within the system, leading to mutual influence between the two oscillators.
From the theorems discussed earlier, we know that the most probable continuous path of the stochastic nearly integrable Hamiltonian system (45) is governed by the deterministic equations in system (46). As $\epsilon\to 0$, the path of system (45) satisfies the large deviation principle, and we can quantify the probability distribution of the system’s deviation from the most probable path using the rate function.
To gain a better understanding of the system’s behavior under different perturbation strengths and to validate our theoretical results, we performed numerical simulations. Specifically, we considered three different perturbation strengths: $\epsilon=0.001$, $\epsilon=0.01$, and $\epsilon=0.1$. The results of these simulations are shown in the figures below.
[…]
At very small perturbations, the system closely resembles an integrable Hamiltonian system, see Fig 2. The invariant tori are well-preserved, and the trajectories in phase space exhibit regular, closed curves. Even with the introduction of deterministic and random perturbations, the system’s trajectory remains largely stable, with minimal impact from the stochastic terms.
Figure 3: The comparison figure of Fig 2 when $\epsilon=0.01$ and $\epsilon=0.1$, respectively.
As the perturbation strength increases, the system’s trajectories begin to change, as shown in Fig 3. While the invariant tori are still present in phase space, the combined effects of deterministic and random perturbations lead to increased complexity in the trajectories. The stability of these trajectories gradually decreases, though they still exhibit quasi-periodic behavior. When the perturbation strength is further increased, the complexity grows significantly, with random perturbations introducing more fluctuations that result in chaotic trajectories. This indicates that when the perturbations are strong enough, the overall topological structure of the system becomes disrupted, ultimately leading to the destruction of the invariant tori.
From the results of these numerical simulations, we can observe that under small stochastic perturbations, the system’s trajectories generally evolve along deterministic paths, and the invariant tori are well-preserved. However, as the perturbation strength increases, the system’s trajectories become progressively more complex. The fluctuations introduced by random perturbations become more pronounced, especially when $\epsilon=0.1$, where the system exhibits stronger chaotic behavior.
Combining theoretical analysis with numerical simulation results, we can draw the conclusion that when the perturbation coefficient is small, the preservation of the invariant torus in the almost integrable stochastic Hamiltonian system can be guaranteed in a probabilistic sense. At this time, the trajectory of the system evolves along the solution of the deterministic Hamiltonian equation, and the probability of deviating from the most likely path decays exponentially. Although random perturbations introduce complexity, as long as the perturbations remain small, the basic structure of the system can still be retained.